A Complete Bibliography of Publications in Synthese, 1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.02

Title word cross-reference

1 [185]. _{2}^{−} [186]. A [186]. n [463].

-sections [185]. -tuples [463].

1 [333].

= [48].

Absolute [61]. abstract [184, 185]. Abstracts [33, 43, 56, 68, 146, 190]. acceptance [150]. according [532]. account [530, 517]. accumulation [115]. accuracy [229, 277, 281]. Achilles [395]. action [238]. actions [20]. activity [530]. added [11]. addition [357]. adequacy [321]. adequate [213, 398]. ado [230]. Adverbials [238]. Adverbs [237, 338]. Afterthoughts [322]. Agassi [264, 263, 265]. age [66]. agnoiology [298]. Alan [167]. Alberto [223]. Algebraic [355]. alive [505]. Allan [412, 413]. alternative [85]. amalgamation [573]. ambiguity [222, 223]. Anadic [341]. analogy [532, 417, 63]. analysis [14, 389, 459, 508, 451, 309, 85]. Analytic [67, 234, 28]. analytic-synthetic [234]. Analytic/ [28]. analytical [308]. analyticity [133]. Anderson [167]. Announcement [145, 191, 168, 169, 211, 312, 334, 345, 390, 432, 461, 501, 572, 324]. Announcements [410, 534]. Anomalies [479, 392]. Anselm [506]. anti [516]. anti-realism [516]. Antinomian [64]. Antinomien [64]. antinomy [87]. any [421]. appeal [110]. Appearance [228]. Application [451]. Applications [306]. applied [183, 381, 446]. Appreciation [395]. approach [288, 249, 163, 488, 151]. approximate [289]. Approximative [475]. archaeology [25]. argument [414, 560, 454]. Aristotelian [218]. Aristotle [216, 213, 215]. Arne [80]. Arthur [551]. article [518]. Artificial [480]. aspects [492, 558]. assert [161]. assumptions [452]. Athena [523]. atomism [293]. attack [83]. attributes [444]. Austrian [431]. Axiomatical [174]. Axiomatization [487, 173].

backward [100]. bargain [455]. based [150]. basis [193, 252]. Bayesian [414, 272, 426, 153]. be [11, 494, 88, 397]. before [504]. behavioral [355]. Belief [193, 402, 282, 494]. beliefs [484]. believes [12]. Believing [9, 12, 159]. between [415, 472, 417]. bibliography [323]. bifurcation [430]. biological [533, 446]. biology [348]. Birnbaum [412, 413, 416]. bisection [50]. bits [313]. black [558]. body [44, 441]. Bohr [5]. Book [232, 233]. Books [69, 81, 93, 143, 155, 188, 311, 300, 344, 409, 527]. Boolean [552, 368]. Boolean-valued [368]. Boston [92, 109]. Brain [50, 49]. Brand [162]. British [56]. Brouwer [179]. Buchdahl [62]. built [131]. Bulletin [333]. bundle [568].

C. [532, 533]. calculus [525, 250]. call [191]. called [255, 139]. Can [494, 88]. can’t [481]. Carathéodory [556]. Carnap [126, 129, 132, 128, 130, 131]. case [224, 375, 508, 539, 526]. cases [336]. Catastrophe [305, 446]. catastrophes [492]. catch [116]. categories [365]. category [288, 294]. category-theoretic [288]. Causal [361, 95, 392, 96, 97, 525, 360, 397, 151, 161, 98, 398]. causation [162, 14, 398]. cause [51]. causes [383]. century [510, 508]. chance [420]. chances [386]. Change [510, 317, 325, 282, 72, 328]. choice [429]. Chomsky [339]. Chr. [67]. chronic [262]. class [342, 483]. classes [375, 343, 433]. classical [2, 8, 3]. classification [202]. classifications [347]. classificatory [542]. code [498]. Coffa [223]. cognition [497]. cognitive [406]. coherence [286]. colloquium [92, 109]. combinatorial [249]. commanded [494]. Comment [279, 203, 218, 199, 204, 194, 198, 278]. commentary [49]. Comments [272, 280, 265, 273, 274, 268, 264, 275, 476, 223, 537, 263, 269, 459, 217]. commitment [239, 295]. Communications [91, 332]. communities [538]. Comparative [112, 542]. comparing [326]. Competence [339, 496]. competence-performance [496]. Competing [366]. complementarity [221]. completeness [256, 392]. complexity [15]. compromise [269]. Computability [497]. computer [481]. concept [556, 173, 416, 522, 215]. conception [60, 412]. concepts [15, 215]. Conceptual [325, 72]. Concerning [398, 21]. conclusions [290, 418]. conditional [247, 82, 517, 161]. Conditionalization [149]. conditionals [457]. conditions [360]. confidence [424, 416]. confirmation [526]. conflict [307]. consciousness [402, 50]. consensus [438, 298]. consequences [114]. conservation [63]. conservative [133]. Consistency [381]. constituents [472]. constraining [483]. constraints [136]. construct [451]. construction [113]. constructional [229]. constructivism [111, 179]. content [5]. continuity [235]. continuum [450]. Contradiction [510, 140]. contribution [504, 133]. control [306]. controversy [510, 152]. convention [104, 377]. conventionalism [379]. conventionality [105]. Coordinate [107]. Coordinate-free [107]. Copenhagen [4]. coreference [23]. corporeal [48]. Correction [210, 7]. Corrections [16, 34]. correlates [408]. correspondence [260, 5]. Corrigenda [283]. corroboration [267]. cosmology [104]. count [317, 554]. Counterfactuals [490, 388, 399, 282]. Craig [120]. Craigian [226]. credo [340]. criteria [46]. Critical [565, 358, 520, 470, 521, 486, 547, 485, 469, 519, 119, 61]. criticism [414, 111, 395]. Critique [271, 456, 446]. cross [325]. cross-theoretical [325]. crucial [44]. current [24].

David [279, 196, 280, 278]. Davidson [194]. dead [505]. Dean [153]. decision [414, 415, 417]. Decisions [418, 416]. deduction [499]. deductive [290]. Deep [19]. definability [184]. Definition [119, 425, 8]. definitions [229]. degrees [13]. demonstrative [507]. Demonstratives [482]. derived [356]. description [479]. descriptions [464, 22]. design [430]. Determinism [550]. deterministic [151]. difference [369]. Difficulties [454]. Dinosaurs [331]. Diodorus [560]. direction [396]. discrete [253]. discussion [207, 206, 208, 476]. disease [262]. disk [380]. dispositional [384]. Dispositions [383, 387, 385, 386]. distance [472, 473]. distinction [234, 415, 496]. disunity [220]. Division [333, 91, 332]. do [29]. doctrine [420]. does [369, 161]. Donald [194]. Dorothea [219]. double [567]. Dummett [199, 228, 563, 198]. D’un [310]. dynamical [328]. dynamics [488].

Earman [98]. economics [306, 308, 309]. editor [303, 170, 428]. Editorial [1, 57, 79, 125, 157, 212, 411, 261, 94]. egocentric [240]. Ehics [443]. Einstein [257]. Elementary [186]. Empirical [248, 158, 452, 163, 122]. Empiricism [74, 111, 330]. Empiricist [164, 126]. energy [557]. English [80, 341, 504]. entailment [291]. entanglements [376]. epistemic [226, 227]. Epistemological [77, 492, 5]. epistemology [114, 456, 388]. equivalent [357]. Errata [302, 346, 422, 445, 495, 511]. Erratum [477, 559, 449]. Errors [259]. erudition [569]. espace [245]. Essai [245]. essay [420, 245]. Essentialism [235]. estimates [424]. Ethics [441, 442, 395]. Euclidean [406, 405]. evaluation [348]. event [48]. events [237, 48]. evidence [464, 412, 416, 467]. evolution [15, 545, 533]. evolutionary [456]. Exact [424]. example [545]. exchangeability [427]. excuse [553]. existence [407, 87]. existential [407]. expectation [297]. experience [359, 63]. experiences [47]. experimental [348]. experiments [255, 532]. explained [88]. Explanation [349, 214, 325, 403, 224, 4, 147, 387]. explanations [113, 213, 491]. Explanatory [110]. Explicativity [267]. exploding [539]. exposition [462]. extended [226]. extensional [229]. extensions [186, 133].

fact [475, 161]. fact-statements [475]. fallacy [369]. Fallibilism [531]. falsification [271]. falsificationism [73]. feels [481]. Feldman [458]. Feyerabend [83]. Final [459]. Fine [551]. Finetti [427]. Finite [427]. First [207, 63]. flow [101]. Fodor [468]. forcing [368]. Foreword [171]. form [216, 540, 19]. formal [465]. formalization [173]. forms [427, 364]. formulation [99, 517]. formulations [121]. foundational [172]. Foundations [6, 255, 362, 359, 549, 421, 183, 509, 460]. fourteenth [510, 508]. fourteenth-century [510, 508]. frame [40]. Frede [219]. free [107]. Frege [515]. French [245, 310]. frequencies [554, 374, 373]. frequency [425]. Frequentist [419]. Freudenthal [440]. Function [64]. functional [202]. functionalism [478]. functionally [88]. functions [282]. Fund [167]. Funktion [64]. Fuzzy [347, 289, 288].

G. [67]. games [564]. Gathering [467]. Geach [563]. General [206, 37, 207, 208, 175]. Generalization [221, 130, 450]. generalizations [225]. generalized [53]. generate [435]. Generative [336, 394]. genetic [498]. genus [216]. geometries [95]. geometry [105, 380, 378]. German [80, 64]. Giere [273]. Global [309]. god [182]. goo [217]. Good [268, 269]. Goodman [538]. grammar [51, 482, 339, 138, 466]. grammars [448, 39]. grammatical [512, 21, 137]. Grene [217]. Grosseteste [507]. grounds [383]. Grünbaum [105]. Guest [303].

Habermas [135]. hand [491]. Hans [378, 382, 371]. heel [395]. Hempel [222, 223]. hermeneutic [135]. Hidden [555, 259]. hierarchy [515]. High [557]. High-energy [557]. Hintikka [218]. historico [328]. historico-dynamical [328]. History [333, 91, 332, 528, 569]. holes [558]. homogeneous [433]. horses [331]. human [498]. hypotheses [267, 75, 434, 450]. hypothesis [349, 220, 137]. hysteresis [352].

I. [268, 269]. idea [175]. ideal [532]. Ideals [52]. identifying [22]. identity [139, 235]. illness [262]. immediate [47]. imperative [121]. implications [393]. importance [569]. incompleteness [258]. inconsistencies [82]. inconsistency [478]. Indefinite [148]. independence [434, 573, 462].

Indeterminacy [567, 493, 86, 197, 26]. indeterminism [245, 147]. indéterminisme [245]. indexicals [482]. individual [429]. Inducibility [227]. Induction [72, 127, 112, 507, 76]. inductions [452]. Inductive [119, 82, 151, 30, 75, 129, 128, 130, 450, 451]. Inductivism [71, 73]. inexact [13]. infantile [262]. Inference [372, 27, 414, 75, 415, 417]. inferences [290]. inferring [159]. Informal [173, 462]. information [153]. ingredients [218]. innateness [467, 137]. inquiry [166]. instabilities [430]. instrumentalism [120]. intelligence [480]. intensional [38]. International [528, 91, 332, 307, 333]. Interpolation [184]. interpretation [30, 195, 163, 4, 551, 122, 555]. Interpreting [506]. intertheory [327]. introducing [384]. Introduction [243, 411, 303, 428, 261, 94]. intuitionism [564]. Intuitionistic [176]. invalidity [339]. Invisible [491]. Invisible-hand [491]. isomorphism [229]. issue [573]. issues [44].

J [268, 223, 269]. Jamison [153]. Jeffrey [487, 272, 274, 275]. Joseph [265, 264, 263]. Journal [56]. Journals [70, 144, 156, 189, 301, 311, 344, 409, 527].

Kant [59, 60, 365, 67, 58, 61, 99, 65, 63]. Kantian [64]. Kantischen [64]. Karttunen [517]. Katz [340, 140]. Kaufman [227]. kinds [295]. know [10]. Knowing [8, 12, 159, 11, 160, 85, 10]. knowledge [231, 359, 8, 115]. knows [12, 11, 8, 9]. Kochen [259].

Laddaga [573]. Lakatos [566]. Language [564, 49, 135, 513, 468, 464, 17, 35, 512, 394, 88, 498]. languages [80, 335, 39, 483]. law [357]. lawlikeness [60]. Laws [399, 475, 82, 99]. learning [468]. lectures [396]. Lehrer [438]. Leibniz [67]. Lemmon [160]. Levy [186]. Lewis [196]. lifetime [58]. light [106]. Lindley [414, 454]. linguistic [464, 24, 467, 447]. Linguistics [41]. list [80]. locality [552, 550]. Logic [333, 67, 166, 91, 332, 362, 293, 159, 284, 569, 248, 245, 453, 341, 249, 247, 391, 2, 129, 128, 130, 41, 38, 382, 21, 200, 504, 489, 451, 500, 503, 138, 292, 151, 289]. Logical [293, 183, 452, 540, 19, 126, 364, 442, 150, 338]. logically [246]. logician [506]. logico [464]. logico-linguistic [464]. logics [513, 184]. logique [245]. lower [435].

machine [297]. Macroevolution [499]. macroscopic [254]. magnitudes [552]. make [481, 369]. Making [275]. man [505]. Many [366, 294]. many-valued [294]. March [68]. Mass [317, 314, 322, 323]. master [560]. Materialism [46]. Mathematical [304, 549, 307]. mathematics [172, 566, 176, 183, 381]. Matter [217, 216, 320, 315, 215, 65]. maxim [329]. Maximum [308]. Meaning [136, 202, 529, 193]. meaningfulness [158]. means [474]. measurement [356]. mechanical [251]. mechanics [255, 53, 549, 252, 248, 391, 545, 382, 550, 3, 551, 258]. mechanistic [244]. medieval [503]. Memorial [167]. Mental [48, 363, 530, 364, 46, 496]. meta [330, 395]. meta-ethics [395]. meta-philosophy [330]. metaphor [30]. Metaphysical [393]. metascience [489]. metastatistics [420]. method [275]. Methodological [24, 112]. Methodology [333, 91, 332, 431]. methods [450]. Michael [228, 199, 198]. microreductions [123]. might [397]. Miller [279, 280, 278]. Mind [441, 44, 45]. mind-body [44]. Minkowski [96]. mistakes [294]. modal [293, 26, 87, 542, 555]. Modalities [399, 151]. Modality [532, 526]. Model [122, 253, 186, 436, 394, 163, 310, 164]. modèle [310]. modèles [310]. Models [429, 304, 95, 307, 557, 310]. modern [215]. modifiers [21]. monism [44]. Montague [482, 464, 466]. Moravcsik [214]. motion [61, 99]. Much [230]. must [11].

Naess [80]. Naïveté [553]. names [465, 22]. naming [404]. Natural [52, 513, 17, 35, 41, 335, 508, 39, 509]. nature [331, 176]. necessity [524, 531]. Needed [328]. nets [545]. Newcomb [490]. Newtonian [545]. Neyman [414]. No [68, 333, 255, 553, 259, 535]. nominalism [293]. nomological [400]. non [272]. non-Bayesian [272]. Noncompensatory [353]. Nondirected [106]. Normative [460]. Nostalgia [537]. notation [394]. note [1, 57, 79, 125, 157, 212, 255, 352, 420, 153]. Notes [127, 97]. notice [565, 358, 520, 521, 486, 547, 485, 519]. Notices [32, 470, 469]. notion [244]. notions [26]. nouns [317, 314]. numbers [178]. Numerical [348].

Objections [541]. Objectively [433]. Objects [484, 402, 401]. observable [260]. observation [254, 83, 149]. observations [253]. Obstacles [115]. obtain [11]. occasion [484]. Ockham [504]. odds [267]. Old [362, 489]. one [12, 11, 8, 10, 9]. Ontological [239, 295, 557]. ontology [525, 26, 321]. Opacity [23]. open [108]. operational [121]. Operationalism [118]. operationally [384]. Operator [260]. Operator-observable [260]. optimal [306]. optimization [430]. Order [52]. Ordinary [364, 384, 394, 537, 535]. ought [29]. ought-to-do [29]. overview [443].

pain [481]. Palter [62]. paper [206, 218, 153, 217]. papers [191, 62, 537]. parable [25]. Paradox [231, 285, 490, 561, 257]. Part [442]. particles [253]. particulars [240, 239]. partition [182]. partitioning [342]. partitions [343]. pattern [365]. peanut [369]. Pearson [414]. Peirce [529, 530, 532, 533]. Perception [361, 402, 540, 362, 403, 407, 360, 539]. Perceptual [541, 401, 408]. perfect [538]. performance [496, 339]. Petitio [296]. phenomena [181]. phenomenology [116]. philosophical [80, 558]. Philosophy [56, 333, 528, 330, 92, 109, 91, 332, 135, 60, 566, 569, 508, 496, 61, 371, 108, 102, 33, 43, 68, 146, 190]. Physical [426, 402, 552, 355]. Physics [6, 4, 63]. Piagetian [480]. pictures [364]. pieces [313]. plurality [314]. Podolsky [257]. Poincaré’s [102]. point [337, 236, 338]. Popper [456, 272, 282, 271, 351]. Popperianism [330]. positivism [493]. possibility [161]. Possible [544]. posteriori [509]. Postscript [209]. power [110]. practice [372, 384]. Pragmatics [38, 42, 514]. predicates [384, 563, 543, 286]. Predication [320]. predictions [277, 281]. Preface [242, 502, 370, 192, 170]. preference [13, 429]. preferences [13, 353]. premises [290]. preservation [226]. presupposition [517]. primer [554]. principii [296]. principle [455, 5]. principles [177, 308, 63]. Prior [560]. probabilism [71]. Probabilistic [39, 525, 252, 398]. probabilitie [375]. probabilities [53, 374, 435]. Probability [253, 271, 573, 447, 522, 425, 437, 127, 251, 148, 250, 77, 150, 419, 3, 426, 542]. problem [365, 420, 67, 408, 133]. Problems [570, 112, 21, 76, 108]. process [499]. Professor [440]. program [59, 36, 271, 226]. progress [570, 533]. Progressions [307]. Projectible [543]. pronouns [23]. proof [259, 175]. proofs [348, 381]. propensities [526]. Propensity [524, 224]. Proper [465, 22]. proposal [438]. Propositional [508]. psychological [403, 339]. Publications [413].

qualitative [357]. quantification [319]. quantifier [291]. quantifying [567]. quantitative [5, 158]. quantities [316]. Quantity [319]. Quantum [553, 2, 3, 255, 53, 251, 549, 256, 252, 554, 248, 249, 393, 247, 391, 250, 392, 382, 550, 246, 551, 258, 555]. quantum-logically [246]. quantum-mechanical [251]. quark [557]. Quine [196, 201, 197, 567, 141].

Radical [195]. ramification [245]. Ramsey [120]. Ramsey-sentences [120]. random [435]. Randomness [425, 434]. Rational [282, 298]. rationality [74]. ravens [350]. reactive [500]. Realism [165, 553, 330, 387, 516]. realist [164]. realists [554]. reality [351]. realization [45]. reasoning [30, 289]. received [69, 81, 70, 93, 143, 155, 144, 156, 188, 189, 311, 300, 301, 344, 409, 527]. recognition [365]. reconcilable [73]. reconsidered [458]. recursive [512]. reducibility [177]. reduction [329]. Reference [287, 375, 562, 236, 433, 27, 40]. reflection [448]. Reflections [54, 24]. regard [384]. regions [424]. Regulative [64]. Regulativen [64]. Reichenbach [379, 375, 376, 391, 400, 378, 377, 382, 371, 395]. Rejoinder [227]. relation [215]. Relational [75, 54]. relations [96, 327, 435, 137]. Relative [373, 267]. relativistic [357]. Relativity [181, 380, 107, 378]. relevance [453]. remarks [314]. replacement [329]. Replies [266, 196, 270, 276]. Reply [201, 219, 440, 141, 205, 89, 573, 281]. Reports [47]. representation [355]. representations [552, 364, 363]. research [172, 67, 112, 503]. respective [2]. Response [458]. results [249, 462]. Resurrecting [350]. retrospect [132]. Review [548, 439, 571, 518, 31, 134, 117, 187, 546]. reviewed [10]. Reviews [241, 90, 142, 55, 124, 154, 299, 78, 232, 233]. revised [252]. revisited [471]. Richard [272, 274, 275]. Riemannian [406]. right [397, 292]. rise [172]. Robert [507]. role [177, 30]. roles [18, 2]. Ronald [273]. Rosen [257]. Ross [167]. rotating [380]. Rudolf [126]. rule [150]. Russell [329].

S [532, 533]. Savage [414, 454]. Schematism [365]. Science [333, 528, 92, 109, 91, 332, 60, 325, 569, 220, 330, 82, 507, 509, 310, 460, 68, 33, 43, 56, 146, 190, 218]. sciences [352, 220, 115, 355, 446]. scientific [166, 114, 120, 384, 163, 533, 147, 122]. Second [208]. Secondary [229]. Secondary-extensional [229]. sections [185]. Seeing [404, 84]. seem [161]. Selected [80]. Self [562, 479, 27]. self-description [479]. Self-reference [562, 27]. Sellars [206, 203, 204]. semantic [28]. Semantical [25, 240, 133]. Semantics [437, 17, 35, 29, 525, 26, 132, 336, 37, 294, 366, 164]. semiotic [465]. sense [275, 48]. Sensory [359]. sentences [540, 238, 120, 83, 512]. session [207, 208]. set [181, 174, 462, 368, 512]. set-theoretical [174]. Sets [180]. settheoretical [182]. Shoenfield [463]. short [474]. signals [106]. significance [197, 152]. Single [526, 224, 375]. Single-case [526]. singular [563]. size [66]. sloppy [139]. so-called [255, 139]. Social [298, 352, 115, 112, 446]. sociological [114, 113, 112]. Sociology [116]. solution [561]. Some [492, 314, 384, 44, 182, 249, 336, 21, 108, 313, 558, 114]. sorites [369]. space [406, 96, 95, 54, 245, 103, 61, 108, 405, 102]. space-time [96, 95, 103]. speaker [337]. speakers [18]. Special [220, 380]. species [216]. Specker [259]. Spinoza [441, 182, 443, 442, 444]. Spinozistic [343, 342]. stand [234]. standard [567]. statements [148, 400, 475, 158]. stating [420]. Statistical [4, 147, 412, 436, 416]. statistics [421, 419, 426]. status [557, 65]. Stochastic [258]. stone [561]. stratified [568]. Strawson [563, 516]. Stray [354]. structuralist [488]. Structure [228, 402, 251, 182, 19, 103, 442, 106]. Studies [569]. study [58, 508]. Stuff [318]. Subjectivism [262]. Subjects [18]. Subjunctive [225, 457]. Subjunctives [386]. suggested [182]. super [366]. super-truth [366]. Superposition [254]. Support [423]. surrender [116]. surrender-and-catch [116]. survey [119]. Swain [162]. symbolic [394, 138]. symmetric [96]. Symmetry [327]. Syntax [525, 138, 36]. synthetic [234, 28]. system [129, 395]. Systematic [165]. systematization [226, 119, 227]. systems [288, 251, 500, 133].

taxonomy [216]. temperature [556]. Temporally [96]. temps [245]. Tense [513, 514]. terms [317, 563, 322, 323]. test [336, 152]. tests [424]. Their [2, 121]. theorem [120, 427]. theorems [182, 474]. theoretic [288, 163]. theoretical [325, 174, 522]. Theories [315, 121, 114, 326, 436, 475, 545, 163, 544, 110, 374, 122, 488]. Theory [361, 545, 473, 181, 354, 240, 414, 480, 492, 372, 97, 224, 54, 256, 13, 568, 393, 380, 436, 306, 400, 392, 244, 462, 28, 368, 356, 175, 24, 397, 3, 305, 446, 430, 457, 122, 164, 65, 328, 98, 555, 398]. Theory-distance [473]. Theory-nets [545]. There [535, 421]. thesis [197]. things [318, 234, 535]. think [246]. thinking [174]. Three [76]. time [96, 95, 97, 245, 103, 106, 397, 108, 100, 101, 396, 98, 398]. Topology [104]. Toulmin [52]. tradition [272]. transformational [138, 483]. translating [394, 200]. translation [86]. transmitted [290]. travelling [100]. Treatments [466]. Troubles [20]. true [466]. Truth [291, 284, 475, 173, 517, 86, 366]. truth-conditional [517]. Truthlikeness [476, 278]. tuples [463]. Two [396]. Tyche [523]. type [321].

unboundedly [500]. uncertainties [290]. underlying [114, 464]. Unger [537]. Unified [356]. Uniform [123]. Union [333, 91, 332, 528]. unity [325, 50]. universal [448, 342, 450, 335]. universals [408]. upper [435]. usage [153]. usefulness [316]. using [435]. utilities [487, 418]. Utility [13].

V [196, 201]. vague [291, 286]. Vagueness [284, 367, 366, 287]. validity [348, 452, 562, 451]. valued [368, 294]. values [366]. variable [259]. variables [555]. variation [447]. Varieties [542]. velocities [357]. Verisimilitude [279, 474, 471, 473]. versus [406, 366]. view [337, 338, 328]. views [176]. Visual [406, 403, 405]. vogue [367]. Vol [68]. vs [164].

W [67]. W. [196, 201]. Wang [285]. Warranting [459, 458, 389]. Ways [331, 384]. Wheeler [537]. Where [234]. Which [292, 536]. Wilfrid [203, 204]. within [226]. without [231]. Wolff [67]. word [466]. work [323]. working [220]. world [288, 66, 304, 385, 131]. writings [80].

yes [255].

Zur [64].

References

Anonymous:1970:EN

[1] Anonymous. Editorial note. Synthese, 21(1):1, March 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

     link.springer.com/article/10.1007/BF00414185.

Heelan:1970:QCL

[2] Patrick Heelan. Quantum and classical logic: Their respective roles. Synthese, 21(1):2–33, March 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

     article/10.1007/BF00414186.

Sneed:1970:QMC

[3] Joseph D. Sneed. Quantum mechanics and classical probability theory. Synthese, 21(1):34–64, March 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

     article/10.1007/BF00414187.

Schlegel:1970:SEP

[4] Richard Schlegel. Statistical explanation in physics: The Copenhagen interpretation. Synthese, 21(1):65–82, March 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00414188.

Komar:1970:QEC

[5] Arthur Komar. The quantitative epistemological content of Bohr’s Correspondence Principle. Synthese, 21(1):83–92, March 1970. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00414189.

Freudenthal:1970:WAF

[6] Hans Freudenthal. What about foundations of physics. Synthese, 21(1): 93–106, March 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00414190.

Popper:1970:C

[7] K. R. Popper. Correction. Synthese, 21(1):107, March 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00414191;   http://link.
springer.com/content/pdf/10.1007/BF00414191.pdf.

Hilpinen:1970:KOK

[8] Risto Hilpinen. Knowing that one knows and the classical definition of knowledge. Synthese, 21(2):109–132, June 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413541.

Lehrer:1970:BOK

[9] Keith Lehrer. Believing that one knows. Synthese, 21(2):133–140, June 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413542.

Hintikka:1970:KOK

[10] Jaakko Hintikka. ‘Knowing that one knows’ reviewed. Synthese, 21(2): 141–162, June 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413543.

Ginet:1970:WMA

[11] Carl Ginet. What must be added to knowing to obtain knowing that one knows? Synthese, 21(2):163–186, June 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413544.

Castaneda:1970:KBO

[12] Hector-Neri Castañeda. On knowing (or believing) that one knows (or believes). Synthese, 21(2):187–203, June 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413545.

Fishburn:1970:UTI

[13] Peter C. Fishburn. Utility theory with inexact preferences and degrees of preference. Synthese, 21(2):204–221, June 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413546.

Brand:1970:AC

[14] Myles Brand and Marshall Swain. On the analysis of causation. Synthese, 21(2):222–227, June 1970. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413547.

Bronowski:1970:NCE

[15] J. Bronowski. New concepts in the evolution of complexity. Synthese, 21(2):228–246, June 1970. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413548.

Anonymous:1970:Ca

[16] Anonymous. Corrections. Synthese, 21(2):247, June 1970. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413549;   http://link.
springer.com/content/pdf/10.1007/BF00413549.pdf.

Davidson:1970:SNLa

[17] Donald Davidson and Gilbert Harman. Semantics of natural language. Synthese, 21(3–4):249–250, October 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484799.

Fillmore:1970:SSR

[18] Charles J. Fillmore. Subjects, speakers, and roles. Synthese, 21(3–4):251– 274, October 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484800.

Harman:1970:DSL

[19] Gilbert Harman. Deep structure as logical form. Synthese, 21(3–4):275– 297, October 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484801.

Fodor:1970:TAA

[20] J. A. Fodor. Troubles about actions. Synthese, 21(3–4):298–319, October 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00484802.

Parsons:1970:SPC

[21] Terence Parsons. Some problems concerning the logic of grammatical modifiers. Synthese, 21(3–4):320–334, October 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00484803.

Donnellan:1970:PNI

[22] Keith S. Donnellan. Proper names and identifying descriptions. Synthese, 21(3–4):335–358, October 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484804.

Partee:1970:OCP

[23] Barbara Hall Partee. Opacity, coreference, and pronouns. Synthese, 21 (3–4):359–385, October 1970. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484805.

Quine:1970:MRC

[24] W. V. Quine. Methodological reflections on current linguistic theory. Synthese, 21(3–4):386–398, October 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484806. See reply [141].

Scott:1970:SAP

[25] Dana Scott. Semantical archaeology: A parable. Synthese, 21(3–4):399– 407, October 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484807.

Hintikka:1970:SMN

[26] Jaakko Hintikka. The semantics of modal notions and the indeterminacy of ontology. Synthese, 21(3–4):408–424, October 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00484808.

VanFraassen:1970:ISR

[27] Bas C. Van Fraassen. Inference and self-reference. Synthese, 21(3–4):425– 438, October 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484809.

Linsky:1970:ASS

[28] Leonard Linsky. Analytic/ synthetic and semantic theory. Synthese, 21 (3–4):439–448, October 1970. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484810.

Castaneda:1970:SOD

[29] Hector-Neri Castañeda. On the semantics of the ought-to-do. Synthese, 21 (3–4):449–468, October 1970. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484811.

Cohen:1970:RIR

[30] L. Jonathan Cohen and Avishai Margalit. The role of inductive reasoning in the interpretation of metaphor. Synthese, 21(3–4):469–487, October 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00484812.

Juhos:1970:R

[31] B. Juhos. Review. Synthese, 21(3–4):488–493, October 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00484813.

Anonymous:1970:N

[32] Anonymous. Notices. Synthese, 21(3–4):494–495, October 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00484814.

Anonymous:1970:APSa

[33] Anonymous. Abstracts from Philosophy of Science. Synthese, 21(3–4): 496–501, October 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484815.

Anonymous:1970:Cb

[34] Anonymous. Corrections. Synthese, 21(3–4):502, October 1970. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00484816;   http:
//link.springer.com/content/pdf/10.1007/BF00484816.pdf.

Davidson:1970:SNLb

[35] Donald Davidson and Gilbert Harman. Semantics of natural language. Synthese, 22(1–2):1–2, December 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413596.

Geach:1970:PS

[36] P. T. Geach. A program for syntax. Synthese, 22(1–2):3–17, December 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413597.

Lewis:1970:GS

[37] David Lewis. General semantics. Synthese, 22(1–2):18–67, December 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413598.

Montague:1970:PIL

[38] Richard Montague. Pragmatics and intensional logic. Synthese, 22(1– 2):68–94, December 1970. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413599.

Suppes:1970:PGN

[39] Patrick Suppes. Probabilistic grammars for natural languages. Synthese, 22(1–2):95–116, December 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413600.

Wallace:1970:FR

[40] John Wallace. On the frame of reference. Synthese, 22(1–2):117–150, December 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413601.

Lakoff:1970:LNL

[41] George Lakoff. Linguistics and natural logic. Synthese, 22(1–2):151– 271, December 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413602.

Stalnaker:1970:P

[42] Robert C. Stalnaker. Pragmatics. Synthese, 22(1–2):272–289, December 1970. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413603.

Anonymous:1970:APSb

[43] Anonymous. Abstracts from Philosophy of Science. Synthese, 22(1–2): 290–293, December 1970. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413604.

Feigl:1971:SCI

[44] Herbert Feigl. Some crucial issues of mind-body monism. Synthese, 22 (3–4):295–312, May 1971. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413429.

Pribram:1971:RM

[45] Karl H. Pribram. The realization of mind. Synthese, 22(3–4):313–322, May 1971. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413430.

Kim:1971:MCM

[46] Jaegwon Kim. Materialism and the criteria of the mental. Synthese, 22 (3–4):323–345, May 1971. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413431.

Smart:1971:RIE

[47] J. J. C. Smart. Reports of immediate experiences. Synthese, 22(3–4): 346–359, May 1971. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413432.

Odegard:1971:SME

[48] Douglas Odegard. The sense of ‘mental events = corporeal events’. Synthese, 22(3–4):360–368, May 1971. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413433.

Toulmin:1971:BLC

[49] Stephen Toulmin. Brain and language: A commentary. Synthese, 22 (3–4):369–395, May 1971. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413434.

Nagel:1971:BBU

[50] Thomas Nagel. Brain bisection and the unity of consciousness. Synthese, 22(3–4):396–413, May 1971. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413435.

Aronson:1971:GC

[51] Jerrold L. Aronson. On the grammar of ‘cause’. Synthese, 22(3–4): 414–430, May 1971. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413436.

Purtill:1971:TIN

[52] Richard L. Purtill. Toulmin on ideals of natural order. Synthese, 22 (3–4):431–437, May 1971. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413437.

Chari:1971:TGP

[53] C. T. K. Chari. Towards generalized probabilities in quantum mechanics. Synthese, 22(3–4):438–447, May 1971. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413438.

Fine:1971:RRT

[54] Arthur Fine. Reflections on a relational theory of space. Synthese, 22 (3–4):448–481, May 1971. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413439.

Kyburg:1971:R

[55] Henry E. Kyburg, Jr. Reviews. Synthese, 22(3–4):482–493, May 1971. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413440.

Anonymous:1971:ABJ

[56] Anonymous. Abstracts from the British Journal of the Philosophy of Science. Synthese, 22(3–4):494–498, May 1971. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413441.

Anonymous:1971:ENa

[57] Anonymous. Editorial note. Synthese, 23(1):1, August 1971. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00414140.

Martin:1971:LSK

[58] Gottfried Martin. A lifetime’s study of Kant. Synthese, 23(1):2–17, August 1971. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00414141.

Agassi:1971:KP

[59] Joseph Agassi. Kant’s program. Synthese, 23(1):18–23, August 1971. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00414142.

Buchdahl:1971:CLK

[60] Gerd Buchdahl. The conception of lawlikeness in Kant’s philosophy of science. Synthese, 23(1):24–46, August 1971. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00414143.

Palter:1971:ASA

[61] Robert Palter. Absolute space and absolute motion in Kant’s critical philosophy. Synthese, 23(1):47–62, August 1971. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00414144.

Butts:1971:BPP

[62] Robert E. Butts. On Buchdahl’s and Palter’s papers. Synthese, 23(1): 63–74, August 1971. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00414145.

Weizsacker:1971:KFA

[63] Carl Friedrich V. Weizsäcker. Kant’s ‘first analogy of experience’ and conservation principles of physics. Synthese, 23(1):75–95, August 1971. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00414146.

Schafer:1971:RFK

[64] Lothar Schäfer. Zur ‘Regulativen Funktion’ Der Kantischen Antinomien. (German) [Toward the ‘regulative function’ of the Kantian Antinomian]. Synthese, 23(1):96–120, August 1971. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00414147.

Walker:1971:SKT

[65] Ralph C. S. Walker. The status of Kant’s theory of matter. Synthese, 23 (1):121–126, August 1971. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00414148.

Bennett:1971:ASW

[66] Jonathan Bennett. The age and size of the world. Synthese, 23(1):127– 146, August 1971. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00414149.

Lenders:1971:ALG

[67] Winfried Lenders. The analytic logic of G. W. Leibniz and Chr. Wolff: A problem in Kant research. Synthese, 23(1):147–153, August 1971. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00414150.

Anonymous:1971:APS

[68] Anonymous. Abstracts from Philosophy of Science (Vol. 37, No. 1, March, 1970). Synthese, 23(1):154–161, August 1971. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00414151.

Anonymous:1971:BRa

[69] Anonymous. Books received. Synthese, 23(1):162–165, August 1971. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00414152.

Anonymous:1971:JR

[70] Anonymous. Journals received. Synthese, 23(1):166, August 1971. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00414153.

Rosenkrantz:1971:IP

[71] Roger Rosenkrantz. Inductivism and probabilism. Synthese, 23(2–3): 167–205, December 1971. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413626.

Lehrer:1971:ICC

[72] Keith Lehrer. Induction and conceptual change. Synthese, 23(2–3):206– 225, December 1971. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413627.

Vetter:1971:IFR

[73] Herman Vetter. Inductivism and falsificationism reconcilable. Synthese, 23(2–3):226–233, December 1971. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413628.

Creary:1971:ER

[74] Lewis G. Creary. Empiricism and rationality. Synthese, 23(2–3):234– 265, December 1971. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413629.

Hilpinen:1971:RHI

[75] Risto Hilpinen. Relational hypotheses and inductive inference. Synthese, 23(2–3):266–286, December 1971. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413630.

Stemmer:1971:TPI

[76] Nathan Stemmer. Three problems in induction. Synthese, 23(2–3):287– 308, December 1971. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413631.

Kyburg:1971:EP

[77] Henry E. Kyburg, Jr. Epistemological probability. Synthese, 23(2–3): 309–326, December 1971. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413632.

Smokler:1971:R

[78] Howard Smokler, R. Harré, and Risto Hilpinen. Reviews. Synthese, 23(2– 3):327–346, December 1971. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413633.

Anonymous:1971:ENb

[79] Anonymous. Editorial note. Synthese, 23(2–3):347, December 1971. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413634.

Flo:1971:ANS

[80] Olav Flo. Arne Naess: Selected list of his philosophical writings in the English and German languages. 1936–1970. Synthese, 23(2–3):348– 352, December 1971. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413635.

Anonymous:1971:BRb

[81] Anonymous. Books received. Synthese, 23(2–3):353–355, December 1971. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413636.

Nowak:1972:IIC

[82] Stefan Nowak. Inductive inconsistencies and conditional laws of science. Synthese, 23(4):357–373, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00636291.

Hull:1972:FAO

[83] Richard T. Hull. Feyerabend’s attack on observation sentences. Synthese, 23(4):374–399, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00636292.

Howell:1972:S

[84] Robert Howell. Seeing as. Synthese, 23(4):400–422, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00636293.

Swain:1972:AAK

[85] Marshall Swain. An alternative analysis of knowing. Synthese, 23(4): 423–442, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00636294.

Rorty:1972:ITT

[86] Richard Rorty. Indeterminacy of translation and of truth. Synthese, 23 (4):443–462, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00636295.

Niemi:1972:EMA

[87] Gunnar Niemi. On the existence of a modal antinomy. Synthese, 23 (4):463–476, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00636296.

Sampson:1972:CLE

[88] Geoffrey Sampson. Can language be explained functionally? Synthese, 23(4):477–486, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00636297.

Toulmin:1972:R

[89] Stephen Toulmin. Reply. Synthese, 23(4):487–490, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00636298.

Drake:1972:R

[90] F. R. Drake and Trond Berg Eriksen. Reviews. Synthese, 23(4):491– 496, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00636299.

Anonymous:1972:CIU

[91] Anonymous. Communications of the international union of history and philosophy of science division of logic, methodology, and philosophy of science. Synthese, 23(4):497–506, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00636300.

Anonymous:1972:BCPa

[92] Anonymous. Boston colloquium for the philosophy of science (1971–1972). Synthese, 23(4):507–509, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00636301.

Anonymous:1972:BR

[93] Anonymous. Books received. Synthese, 23(4):510, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00636302.

Suppes:1972:EI

[94] Patrick Suppes. Editorial introduction. Synthese, 24(1–2):1, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00540139.

Domotor:1972:CMS

[95] Zoltan Domotor. Causal models and space-time geometries. Synthese, 24(1–2):5–57, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00540140.

Berger:1972:TSC

[96] George Berger. Temporally symmetric causal relations in Minkowski space-time. Synthese, 24(1–2):58–73, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00540141.

Earman:1972:NCT

[97] John Earman. Notes on the causal theory of time. Synthese, 24(1–2): 74–86, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00540142.

vanFraassen:1972:ECT

[98] Bas C. van Fraassen. Earman on the causal theory of time. Synthese, 24(1–2):87–95, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00540143.

Palter:1972:KFL

[99] Robert Palter. Kant’s formulation of the laws of motion. Synthese, 24 (1–2):96–116, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00540144.

Weingard:1972:TBT

[100] Robert Weingard. On travelling backward in time. Synthese, 24(1–2): 117–132, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00540145.

Zwart:1972:FT

[101] P. J. Zwart. The flow of time. Synthese, 24(1–2):133–158, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00540146.

Vuillemin:1972:PPS

[102] Jules Vuillemin. Poincaré’s philosophy of space. Synthese, 24(1–2):161– 179, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00540147.

Harrison:1972:SST

[103] Craig Harrison. On the structure of space-time. Synthese, 24(1–2):180– 194, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00540148.

Glymour:1972:TCC

[104] Clark Glymour. Topology, cosmology and convention. Synthese, 24(1– 2):195–218, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00540149.

Friedman:1972:GCG

[105] Michael Friedman. Grünbaum on the conventionality of geometry. Synthese, 24(1–2):219–235, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00540150.

Latzer:1972:NLS

[106] Robert W. Latzer. Nondirected light signals and the structure of time. Synthese, 24(1–2):236–280, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00540151.

Hudgin:1972:CFR

[107] Richard H. Hudgin. Coordinate-free relativity. Synthese, 24(1–2):281– 297, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00540152.

Suppes:1972:SOP

[108] Patrick Suppes. Some open problems in the philosophy of space and time. Synthese, 24(1–2):298–316, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00540153.

Anonymous:1972:BCPb

[109] Anonymous. Boston colloquium for the philosophy of science (1972–1973). Synthese, 24(1–2):317–320, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00540154.

Rapoport:1972:EPE

[110] Anatol Rapoport. Explanatory power and explanatory appeal of theories. Synthese, 24(3–4):321–342, September 1972. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413651.

Galtung:1972:ECC

[111] Johan Galtung. Empiricism, criticism, constructivism. Synthese, 24(3–4): 343–372, September 1972. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413652.

Nowak:1972:CSR

[112] Stefan Nowak. Comparative social research and methodological problems of sociological induction. Synthese, 24(3–4):373–400, September 1972. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413653.

Cohen:1972:CSE

[113] Bernard P. Cohen. On the construction of sociological explanations. Synthese, 24(3–4):401–409, September 1972. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413654.

Boudon:1972:UES

[114] Raymond Boudon. On the underlying epistemology of some sociological theories and on its scientific consequences. Synthese, 24(3–4):410– 430, September 1972. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413655.

Lee:1972:OAK

[115] Hans E. Lee. Obstacles to the accumulation of knowledge in the social sciences. Synthese, 24(3–4):431–438, September 1972. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413656.

Wolff:1972:SPS

[116] Kurt H. Wolff. Sociology, phenomenology, and surrender-and-catch. Synthese, 24(3–4):439–471, September 1972. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413657.

Rosenkrantz:1972:R

[117] Roger D. Rosenkrantz. Review. Synthese, 24(3–4):472–474, September 1972. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413658.

Gillies:1972:O

[118] D. A. Gillies. Operationalism. Synthese, 25(1–2):1–24, March 1972. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00484997.

Niiniluoto:1972:ISD

[119] Ilkka Niiniluoto. Inductive systematization: Definition and a critical survey. Synthese, 25(1–2):25–81, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00484998.

Cornman:1972:CTR

[120] James W. Cornman. Craig’s theorem, Ramsey-sentences, and scientific instrumentalism. Synthese, 25(1–2):82–128, March 1972. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00484999.

Suppe:1972:TTF

[121] Frederick Suppe. Theories, their formulations, and the operational imperative. Synthese, 25(1–2):129–164, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485000.

Tuomela:1972:MTE

[122] Raimo Tuomela. Model theory and empirical interpretation of scientific theories. Synthese, 25(1–2):165–175, March 1972. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485001.

Causey:1972:UM

[123] Robert L. Causey. Uniform microreductions. Synthese, 25(1–2):176– 218, March 1972. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485002.

Martin:1972:R

[124] Michael Martin, Robert L. Causey, Ernest W. Adams, Peter Achinstein, and Peter Caws. Reviews. Synthese, 25(1–2):219–253, March 1972. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485003.

Anonymous:1973:EN

[125] Anonymous. Editorial note. Synthese, 25(3–4):255, April 1973. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00499680.

Hempel:1973:RCL

[126] Carl G. Hempel. Rudolf Carnap, logical empiricist. Synthese, 25(3–4): 256–268, April 1973. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00499681.

Carnap:1973:NPI

[127] Rudolf Carnap. Notes on probability and induction. Synthese, 25(3–4): 269–298, April 1973. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00499682.

Jeffrey:1973:CIL

[128] Richard C. Jeffrey. Carnap’s inductive logic. Synthese, 25(3–4):299– 306, April 1973. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00499683.

Hilpinen:1973:CNS

[129] Risto Hilpinen. Carnap’s new system of inductive logic. Synthese, 25 (3–4):307–333, April 1973. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00499684.

Kuipers:1973:GCI

[130] Theo A. F. Kuipers. A generalization of Carnap’s inductive logic. Synthese, 25(3–4):334–336, April 1973. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00499685.

Wedberg:1973:HCB

[131] Anders Wedberg. How Carnap built the world in 1928. Synthese, 25 (3–4):337–371, April 1973. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00499686.

Hintikka:1973:CSR

[132] Jaakko Hintikka. Carnap’s semantics in retrospect. Synthese, 25(3–4): 372–397, April 1973. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00499687.

Williams:1973:CES

[133] P. M. Williams. On the conservative extensions of semantical systems: A contribution to the problem of analyticity. Synthese, 25(3–4):398– 416, April 1973. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00499688.

Niiniluoto:1973:R

[134] Ilkka Niiniluoto. Review. Synthese, 25(3–4):417–436, April 1973. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00499689.

Bar-Hillel:1973:HHP

[135] Y. Bar-Hillel. On Habermas’ hermeneutic philosophy of language. Synthese, 26(1):1–12, October 1973. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869753.

Schnelle:1973:MC

[136] H. Schnelle. Meaning constraints. Synthese, 26(1):13–37, October 1973. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869754.

Wasow:1973:IHG

[137] Thomas Wasow. The innateness hypothesis and grammatical relations. Synthese, 26(1):38–56, October 1973. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869755.

Stenius:1973:SSL

[138] Erik Stenius. Syntax of symbolic logic and transformational grammar. Synthese, 26(1):57–80, October 1973. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869756.

Dahl:1973:CSI

[139] Östen Dahl. On so-called ‘sloppy identity’. Synthese, 26(1):81–112, October 1973. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869757.

Davis:1973:KC

[140] Steven Davis. Katz on contradiction. Synthese, 26(1):113–121, October 1973. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00869758.

Sellars:1973:RQ

[141] Wilfrid Sellars. Reply to Quine. Synthese, 26(1):122–145, October 1973. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869759. See [24].

Hanson:1973:R

[142] William H. Hanson, Gilbert Harman, N. L. Wilson, M. J. Cresswell, Storrs McCall, and Margaret D. Wilson. Reviews. Synthese, 26(1):146– 178, October 1973. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869760.

Anonymous:1973:BRa

[143] Anonymous. Books received. Synthese, 26(1):179–186, October 1973. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869761.

Anonymous:1973:JRa

[144] Anonymous. Journals received. Synthese, 26(1):187, October 1973. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869762.

Anonymous:1973:A

[145] Anonymous. Announcement. Synthese, 26(1):188, October 1973. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869763.

Anonymous:1973:APS

[146] Anonymous. Abstracts from Philosophy of Science. Synthese, 26(1):189– 196, October 1973. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869764.

Srzednicki:1973:SIS

[147] Jan Srzednicki. Statistical indeterminism and scientific explanation. Synthese, 26(2):197–204, December 1973. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00873262.

Jackson:1973:IPS

[148] Frank Jackson and Robert Pargetter. Indefinite probability statements. Synthese, 26(2):205–217, December 1973. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00873263.

Teller:1973:CO

[149] Paul Teller. Conditionalization and observation. Synthese, 26(2):218– 258, December 1973. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00873264.

Mortimer:1973:RAB

[150] Halina Mortimer. A rule of acceptance based on logical probability. Synthese, 26(2):259–263, December 1973. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00873265.

Uchii:1973:ILC

[151] Soshichi Uchii. Inductive logic with causal modalities: A deterministic approach. Synthese, 26(2):264–303, December 1973. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00873266.

Rosenkrantz:1973:STC

[152] R. D. Rosenkrantz. The significance test controversy. Synthese, 26(2): 304–321, December 1973. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00873267.

Mustonen:1973:NDJ

[153] Seppo Mustonen. A note on Dean Jamison’s paper ‘Bayesian information usage’. Synthese, 26(2):322–323, December 1973. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00873268.

Martin:1973:R

[154] Michael Martin and Robert C. Coburn. Reviews. Synthese, 26(2):324– 338, December 1973. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00873269.

Anonymous:1973:BRb

[155] Anonymous. Books received. Synthese, 26(2):339–341, December 1973. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00873270.

Anonymous:1973:JRb

[156] Anonymous. Journals received. Synthese, 26(2):342, December 1973. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00873271.

Anonymous:1974:ENa

[157] Anonymous. Editorial note. Synthese, 26(3–4):343, April 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00883098.

Przelecki:1974:EMQ

[158] Marian Przelecki. Empirical meaningfulness of quantitative statements. Synthese, 26(3–4):344–355, April 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00883099.

Eberle:1974:LBK

[159] Rolf A. Eberle. A logic of believing, knowing, and inferring. Synthese, 26 (3–4):356–382, April 1974. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00883100.

Mannison:1974:LK

[160] D. S. Mannison. Lemmon on knowing. Synthese, 26(3–4):383–390, April 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00883101.

White:1974:WDC

[161] Morton White. Why does a causal conditional seem to assert possibility when in fact it does not? Synthese, 26(3–4):391–395, April 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00883102.

Barker:1974:BSC

[162] John A. Barker. Brand and Swain on causation. Synthese, 26(3–4): 396–400, April 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00883103.

Przelecki:1974:MTA

[163] Marian Przelecki. On model theoretic approach to empirical interpretation of scientific theories. Synthese, 26(3–4):401–406, April 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00883104.

Tuomela:1974:EVR

[164] Raimo Tuomela. Empiricist vs. realist semantics and model theory. Synthese, 26(3–4):407–408, April 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00883105.

Hooker:1974:SR

[165] C. A. Hooker. Systematic realism. Synthese, 26(3–4):409–497, April 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00883106.

Agassi:1974:LSI

[166] Joseph Agassi. The logic of scientific inquiry. Synthese, 26(3–4):498– 514, April 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00883107.

Anonymous:1974:ARA

[167] Anonymous. Alan Ross Anderson Memorial Fund. Synthese, 26(3–4): 515, April 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00883108.

Anonymous:1974:Aa

[168] Anonymous. Announcement. Synthese, 26(3–4):516–517, April 1974. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00883109.

Anonymous:1974:Ab

[169] Anonymous. Announcement. Synthese, 26(3–4):518, April 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00883110.

Anonymous:1974:EP

[170] Anonymous. Editor’s preface. Synthese, 27(1–2):1, May/June 1974. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00660884.

Agazzi:1974:F

[171] E. Agazzi. Foreword. Synthese, 27(1–2):3–5, May/June 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00660885.

Agazzi:1974:RFR

[172] Evandro Agazzi. The rise of the foundational research in mathematics. Synthese, 27(1–2):7–26, May/June 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00660886.

Parsons:1974:IAF

[173] Charles Parsons. Informal axiomatization, formalization and the concept of truth. Synthese, 27(1–2):27–47, May/June 1974. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00660887.

Casari:1974:AST

[174] Ettore Casari. Axiomatical and set-theoretical thinking. Synthese, 27(1– 2):49–61, May/June 1974. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00660888.

Prawitz:1974:IGP

[175] Dag Prawitz. On the idea of a general proof theory. Synthese, 27(1– 2):63–77, May/June 1974. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00660889.

Heyting:1974:IVN

[176] Arend Heyting. Intuitionistic views on the nature of mathematics. Synthese, 27(1–2):79–91, May/June 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00660890.

Cellucci:1974:RRP

[177] Carlo Cellucci. On the role of reducibility principles. Synthese, 27(1–2): 93–110, May/June 1974. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00660891.

White:1974:WN

[178] Nicholas P. White. What numbers are. Synthese, 27(1–2):111–124, May/June 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00660892.

Posy:1974:BC

[179] Carl J. Posy. Brouwer’s constructivism. Synthese, 27(1–2):125–159, May/June 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00660893.

Stenius:1974:S

[180] Erik Stenius. Sets. Synthese, 27(1–2):161–188, May/June 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00660894.

Aaberg:1974:RPS

[181] Claes Åberg. Relativity phenomena in set theory. Synthese, 27(1–2): 189–198, May/June 1974. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00660895.

Friedman:1974:SSP

[182] Joel Friedman. Some settheoretical partition theorems suggested by the structure of Spinoza’s god. Synthese, 27(1–2):199–209, May/June 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00660896.

Nalimov:1974:LFA

[183] V. V. Nalimov. Logical foundations of applied mathematics. Synthese, 27(1–2):211–250, May/June 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00660897.

Jensen:1974:IDA

[184] Finn V. Jensen. Interpolation and definability in abstract logics. Synthese, 27(1–2):251–257, May/June 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00660898.

Normann:1974:AS

[185] Dag Normann. On abstract 1-sections. Synthese, 27(1–2):259–263, May/June 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00660899.

Guzicki:1974:EEL

[186] Wojciech Guzicki. Elementary extensions of Levy’s model of A _{2}^{−}. Synthese, 27(1–2):265–270, May/June 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00660900.

Scanlon:1974:R

[187] T. M. Scanlon. Review. Synthese, 27(1–2):271–284, May/June 1974. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00660901.

Anonymous:1974:BR

[188] Anonymous. Books received. Synthese, 27(1–2):285–289, May/June 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00660902.

Anonymous:1974:JR

[189] Anonymous. Journals received. Synthese, 27(1–2):291, May/June 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00660903.

Anonymous:1974:APS

[190] Anonymous. Abstracts from Philosophy of Science. Synthese, 27(1–2): 293–303, May/June 1974. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00660904.

Anonymous:1974:ACP

[191] Anonymous. Announcement and call for papers. Synthese, 27(1–2): 305, May/June 1974. CODEN SYNTAE. ISSN 0039-7857(print) 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00660905.

Troyer:1974:P

[192] John G. Troyer and Samuel C. Wheeler III. Preface. Synthese, 27(3– 4):307, September 1974. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484596.

Davidson:1974:BBM

[193] Donald Davidson. Belief and the basis of meaning. Synthese, 27(3–4): 309–323, September 1974. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484597.

Quine:1974:CDD

[194] W. V. Quine. Comment on Donald Davidson. Synthese, 27(3–4):325– 329, September 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/

BF00484598. See reply [196].

Lewis:1974:RI

[195] David Lewis. Radical interpretation. Synthese, 27(3–4):331–344, September 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/

BF00484599. See reply [196].

Davidson:1974:RDL

[196] Donald Davidson. Replies to David Lewis and W. V. Quine. Synthese, 27(3–4):345–349, September 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484600. See [195, 194].

Dummett:1974:SQI

[197] Michael Dummett. The significance of Quine’s indeterminacy thesis. Synthese, 27(3–4):351–397, September 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484601.

Quine:1974:CMD

[198] W. V. Quine. Comment on Michael Dummett. Synthese, 27(3–4):399, September 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/

BF00484602. See reply [201].

Harman:1974:CMD

[199] Gilbert Harman. Comment on Michael dummett. Synthese, 27(3–4): 401–404, September 1974. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484603.

Parsons:1974:TL

[200] Charles Parsons. On translating logic. Synthese, 27(3–4):405–411, September 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484604.

Dummett:1974:RWV

[201] Michael Dummett. Reply to W. V. Quine. Synthese, 27(3–4):413–416, September 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484605. See [198].

Sellars:1974:MFC

[202] Wilfrid Sellars. Meaning as functional classification. Synthese, 27(3–4): 417–437, September 1974. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.

1007/BF00484606. See [203, 204] and reply [205].

Dennett:1974:CWS

[203] Daniel Dennett. Comment on Wilfrid Sellars. Synthese, 27(3–4):439– 444, September 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484607. See [202].

Putnam:1974:CWS

[204] Hilary Putnam. Comment on Wilfrid Sellars. Synthese, 27(3–4):445– 455, September 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484608. See [202].

Sellars:1974:R

[205] Wilfrid Sellars. Reply. Synthese, 27(3–4):457–466, September 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00484609. See

[202, 203, 204].

Anonymous:1974:GDS

[206] Anonymous. General discussion of Sellars’s paper. Synthese, 27(3–4): 467–470, September 1974. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484610.

Anonymous:1974:FGD

[207] Anonymous. First general discussion session. Synthese, 27(3–4):471– 508, September 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484611.

Anonymous:1974:SGD

[208] Anonymous. Second general discussion session. Synthese, 27(3–4):509– 521, September 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484612.

Dummett:1974:P

[209] Michael Dummett. ‘Postscript’. Synthese, 27(3–4):523–534, September 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00484613.

Anonymous:1974:C

[210] Anonymous. Correction. Synthese, 27(3–4):535–536, September 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00484614;   http:
//link.springer.com/content/pdf/10.1007/BF00484614.pdf.

Anonymous:1974:Ac

[211] Anonymous. Announcement. Synthese, 27(3–4):537, September 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00484615.

Anonymous:1974:ENb

[212] Anonymous. Editorial note. Synthese, 28(1):1, September 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869492.

Moravcsik:1974:AAE

[213] Julius M. E. Moravcsik. Aristotle on adequate explanations. Synthese, 28 (1):3–17, September 1974. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869493.

Bogen:1974:ME

[214] James Bogen. Moravcsik on explanation. Synthese, 28(1):19–25, September 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869494.

Suppes:1974:ACM

[215] Patrick Suppes. Aristotle’s concept of matter and its relation to modern concepts of matter. Synthese, 28(1):27–50, September 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869495.

Anonymous:1974:GSM

[216] Anonymous. Is genus to species as matter to form? Aristotle and taxonomy. Synthese, 28(1):51–69, September 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00869496.

Rorty:1974:MGC

[217] Richard Rorty. Matter as goo: comments on Grene’s paper. Synthese, 28 (1):71–77, September 1974. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869497.

Frede:1974:CHP

[218] Dorothea Frede. Comment on Hintikka’s paper ‘On the ingredients of an Aristotelian science’. Synthese, 28(1):79–89, September 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869498. See reply [219].

Hintikka:1974:RDF

[219] Jaakko Hintikka. Reply to Dorothea Frede. Synthese, 28(1):91–96, September 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869499. See [218].

Fodor:1974:SSD

[220] J. A. Fodor. Special sciences (or: The disunity of science as a working hypothesis). Synthese, 28(2):97–115, October 1974. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485230.

Lindenberg:1974:GC

[221] Siegwart Lindenberg and Paul Oppenheim. Generalization of complementarity. Synthese, 28(2):117–139, October 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485231.

Coffa:1974:HA

[222] J. Alberto Coffa. Hempel’s ambiguity. Synthese, 28(2):141–163, October 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485232. See

comments [223].

Salmon:1974:CHA

[223] Wesley C. Salmon. Comments on ‘Hempel’s ambiguity’ by J. Alberto Coffa. Synthese, 28(2):165–169, October 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485233. See [222].

Fetzer:1974:SCP

[224] James H. Fetzer. A single case propensity theory of explanation. Synthese, 28(2):171–198, October 1974. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485234.

Pollock:1974:SG

[225] John L. Pollock. Subjunctive generalizations. Synthese, 28(2):199– 214, October 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485235.

Kaufman:1974:PES

[226] Steven A. Kaufman. The preservation of epistemic systematization within the extended Craigian program. Synthese, 28(2):215–221, October 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485236.

Niiniluoto:1974:IES

[227] Ilkka Niiniluoto. Inducibility and epistemic systematization: Rejoinder to Kaufman. Synthese, 28(2):223–232, October 1974. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485237.

Farrell:1974:MDS

[228] Robert Farrell. Michael Dummett on the structure of appearance. Synthese, 28(2):233–249, October 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485238.

Lincicome:1974:SEI

[229] David Lincicome. Secondary-extensional isomorphism and the accuracy of constructional definitions. Synthese, 28(2):251–257, October 1974. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485239.

Goodman:1974:MA

[230] Nelson Goodman. Much ado. Synthese, 28(2):259, October 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485240.

Barker:1974:PK

[231] John A. Barker. Paradox without knowledge. Synthese, 28(2):261– 270, October 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485241.

Bogdan:1974:BR

[232] Radu J. Bogdan. Book reviews. Synthese, 28(2):271–277, October 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485242.

Michalos:1974:BR

[233] Alex C. Michalos. Book reviews. Synthese, 28(2):279–281, October 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485243.

Katz:1974:WTN

[234] Jerrold J. Katz. Where things now stand with the analytic-synthetic distinction. Synthese, 28(3–4):283–319, November 1974. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00877579.

Wiggins:1974:ECI

[235] David Wiggins. Essentialism, continuity, and identity. Synthese, 28(3–4): 321–359, November 1974. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00877580.

Richards:1974:PR

[236] Barry Richards. A point of reference. Synthese, 28(3–4):361–454, November 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00877581.

Cresswell:1974:AE

[237] M. J. Cresswell. Adverbs and events. Synthese, 28(3–4):455–481, November 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00877582.

Borowski:1974:AAS

[238] E. J. Borowski. Adverbials in action sentences. Synthese, 28(3–4):483– 512, November 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00877583.

Jubien:1974:OCP

[239] Michael Jubien. Ontological commitment to particulars. Synthese, 28(3– 4):513–531, November 1974. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00877584.

Barnett:1974:NST

[240] Dene Barnett. A new semantical theory of egocentric particulars. Synthese, 28(3–4):533–547, November 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00877585.

Bynum:1974:R

[241] Terrell Ward Bynum, William G. Lycan, and Ronald E. Nusenoff. Reviews. Synthese, 28(3–4):549–559, November 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00877586.

Anonymous:1974:P

[242] Anonymous. Preface. Synthese, 29(1–4):1, December 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00484947.

Suppes:1974:I

[243] Patrick Suppes. Introduction. Synthese, 29(1–4):3–8, December 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00484948.

Kreisel:1974:NMT

[244] G. Kreisel. A notion of mechanistic theory. Synthese, 29(1–4):11–26, December 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484949.

Fraisse:1974:ELI

[245] Roland Fraïssé. Essai sur la logique de l’indéterminisme et la ramification de l’espace–temps. (French) [Essay on the logic of indeterminism and the ramification of space–time]. Synthese, 29(1–4):27–54, December 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00484950.

Putnam:1974:HTQ

[246] Hilary Putnam. How to think quantum-logically. Synthese, 29(1–4):55– 61, December 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484951.

Hardegree:1974:CQL

[247] Gary M. Hardegree. The conditional in quantum logic. Synthese, 29(1– 4):63–80, December 1974. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484952.

Foulis:1974:ELQ

[248] D. J. Foulis and C. H. Randall. Empirical logic and quantum mechanics. Synthese, 29(1–4):81–111, December 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484953.

Greechie:1974:SRC

[249] Richard J. Greechie. Some results from the combinatorial approach to quantum logic. Synthese, 29(1–4):113–127, December 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http:

//link.springer.com/article/10.1007/BF00484954.

Jauch:1974:QPC

[250] J. M. Jauch. The quantum probability calculus. Synthese, 29(1–4):131– 154, December 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484955.

Domotor:1974:PSQ

[251] Zoltan Domotor. The probability structure of quantum-mechanical systems. Synthese, 29(1–4):155–185, December 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00484956.

Fine:1974:TRP

[252] Terrence L. Fine. Towards a revised probabilistic basis for quantum mechanics. Synthese, 29(1–4):187–201, December 1974. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00484957.

Bastin:1974:PDM

[253] Ted Bastin. Probability in a discrete model of particles and observations. Synthese, 29(1–4):203–227, December 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00484958.

Cartwright:1974:SMO

[254] Nancy Delaney Cartwright. Superposition and macroscopic observation. Synthese, 29(1–4):229–242, December 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00484959.

Bjornestad:1974:NCY

[255] Øistein Bjørnestad. A note on the so-called yes–no experiments and the foundations of quantum mechanics. Synthese, 29(1–4):243–253, December 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00484960.

Fine:1974:CQT

[256] Arthur Fine. On the completeness of quantum theory. Synthese, 29(1–4): 257–289, December 1974. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484961.

VanFraassen:1974:EPR

[257] Bas C. Van Fraassen. The Einstein–Podolsky–Rosen paradox. Synthese, 29(1–4):291–309, December 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484962.

Suppes:1974:SIQ

[258] Patrick Suppes and Mario Zanotti. Stochastic incompleteness of quantum mechanics. Synthese, 29(1–4):311–330, December 1974. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00484963.

Latzer:1974:ENH

[259] Robert W. Latzer. Errors in the no hidden variable proof of Kochen and Specker. Synthese, 29(1–4):331–372, December 1974. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00484964.

Ross:1974:OOC

[260] David J. Ross. Operator-observable correspondence. Synthese, 29(1–4): 373–403, December 1974. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484965.

Goode:1975:EI

[261] Terry M. Goode, Barry M. Loewer, Roger D. Rosenkrantz, and John R. Wettersten. Editorial introduction. Synthese, 30(1–2):1, February/March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485288.

Agassi:1975:SID

[262] Joseph Agassi. Subjectivism: From infantile disease to chronic illness. Synthese, 30(1–2):3–14, February/March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485289.

Spielman:1975:CJA

[263] Stephen Spielman. Comments on Joseph agassi. Synthese, 30(1–2):15– 23, February/March 1975. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485290.

Loewer:1975:CJA

[264] Barry Loewer. Comments on Joseph agassi. Synthese, 30(1–2):25–30, February/March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485291.

Good:1975:CJA

[265] I. J. Good. Comments on Joseph Agassi. Synthese, 30(1–2):31, February/March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/

BF00485292. See [262], comments [268, 269], and reply [270].

Agassi:1975:R

[266] Joseph Agassi. Replies. Synthese, 30(1–2):33–38, February/March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485293.

Good:1975:ECR

[267] Irving John Good. Explicativity, corroboration, and the relative odds of hypotheses. Synthese, 30(1–2):39–73, February/March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485294.

Harper:1975:CJG

[268] William L. Harper. Comments on I. J. Good. Synthese, 30(1–2):75– 78, February/March 1975. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485295. See [265].

Wettersten:1975:GCC

[269] John R. Wettersten. Good’s compromise: Comments on I. J. Good. Synthese, 30(1–2):79–82, February/March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485296. See [265].

Good:1975:R

[270] I. J. Good. Replies. Synthese, 30(1–2):83–93, February/March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485297. See

[265, 262, 268, 269].

Jeffrey:1975:PFC

[271] Richard C. Jeffrey. Probability and falsification: Critique of the Popper program. Synthese, 30(1–2):95–117, February/March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485298.

Giere:1975:PNB

[272] Ronald N. Giere. Popper and the non-Bayesian tradition: Comments on Richard Jeffrey. Synthese, 30(1–2):119–132, February/March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485299.

Good:1975:CRG

[273] I. J. Good. Comments on ronald giere. Synthese, 30(1–2):133, February/March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485300.

Goode:1975:CRJ

[274] Terry M. Goode. Comments on Richard Jeffrey. Synthese, 30(1–2):135– 138, February/March 1975. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485301.

Miller:1975:MSM

[275] David Miller. Making sense of method: Comments on Richard Jeffrey. Synthese, 30(1–2):139–147, February/March 1975. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485302.

Jeffrey:1975:R

[276] Richard C. Jeffrey. Replies. Synthese, 30(1–2):149–157, February/March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485303.

Miller:1975:AP

[277] David Miller. The accuracy of predictions. Synthese, 30(1–2):159–191, February/March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/

BF00485304. See comments [278, 279, 280] and reply [281].

Rosenkrantz:1975:TCD

[278] Roger D. Rosenkrantz. Truthlikeness: Comment on David Miller. Synthese, 30(1–2):193–197, February/March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485305. See [277] and reply [281].

Agassi:1975:VCD

[279] Joseph Agassi. Verisimilitude: Comment on David Miller. Synthese, 30 (1–2):199–204, February/March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485306. See [277] and reply [281].

Good:1975:CDM

[280] I. J. Good. Comments on David Miller. Synthese, 30(1–2):205–206, February/March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/

BF00485307. See [277] and reply [281].

Miller:1975:APR

[281] David Miller. The accuracy of predictions: A reply. Synthese, 30(1–2):207– 219, February/March 1975. CODEN SYNTAE. ISSN 0039-7857(print)

1573-0964 (electronic). URL http://link.springer.com/article/10.

1007/BF00485308. See comments [278, 279, 280].

Harper:1975:RBC

[282] William L. Harper. Rational belief change, Popper functions and counterfactuals. Synthese, 30(1–2):221–262, February/March 1975. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485309.

Anonymous:1975:C

[283] Anonymous. Corrigenda. Synthese, 30(1–2):263, February/March 1975. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485310;
https://link.springer.com/content/pdf/10.1007/BF00485310.pdf.

Fine:1975:VTL

[284] Kit Fine. Vagueness, truth and logic. Synthese, 30(3–4):265–300, September 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485047.

Dummett:1975:WP

[285] Michael Dummett. Wang’s paradox. Synthese, 30(3–4):301–324, September 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485048.

Wright:1975:CVP

[286] Crispin Wright. On the coherence of vague predicates. Synthese, 30(3–4): 325–365, September 1975. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485049.

Wheeler:1975:RV

[287] Samuel C. Wheeler III. Reference and vagueness. Synthese, 30(3–4): 367–379, September 1975. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485050.

Arbib:1975:CTA

[288] Michael A. Arbib and Ernest G. Manes. A category-theoretic approach to systems in a fuzzy world. Synthese, 30(3–4):381–406, September 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485051.

Zadeh:1975:FLA

[289] L. A. Zadeh. Fuzzy logic and approximate reasoning. Synthese, 30(3–4): 407–428, September 1975. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485052.

Adams:1975:UTP

[290] Ernest W. Adams and Howard P. Levine. On the uncertainties transmitted from premises to conclusions in deductive inferences. Synthese, 30(3–4): 429–460, September 1975. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485053.

Carlstrom:1975:TEV

[291] Ian F. Carlstrom. Truth and entailment for a vague quantifier. Synthese, 30(3–4):461–495, September 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485054.

Tharp:1975:WLR

[292] Leslie H. Tharp. Which logic is the right logic? Synthese, 31(1):1–21, June 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00869469.

Cocchiarella:1975:LAN

[293] Nino Cocchiarella. Logical atomism, nominalism, and modal logic. Synthese, 31(1):23–62, June 1975. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869470.

Martin:1975:MVS

[294] John Martin. A many-valued semantics for category mistakes. Synthese, 31(1):63–83, June 1975. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869471.

Jubien:1975:OCK

[295] Michael Jubien. Ontological commitment to kinds. Synthese, 31(1): 85–106, June 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869472.

Woods:1975:PP

[296] John Woods and Douglas Walton. Petitio principii. Synthese, 31(1): 107–127, June 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869473.

Nelson:1975:ME

[297] R. J. Nelson. On machine expectation. Synthese, 31(1):129–139, June 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00869474.

Lehrer:1975:SCR

[298] Keith Lehrer. Social consensus and rational agnoiology. Synthese, 31(1): 141–160, June 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869475.

Niiniluoto:1975:R

[299] Ilkka Niiniluoto and Ronald Giere. Reviews. Synthese, 31(1):161– 199, June 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869476.

Anonymous:1975:BR

[300] Anonymous. Books received. Synthese, 31(1):201–204, June 1975. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869477.

Anonymous:1975:JR

[301] Anonymous. Journals received. Synthese, 31(1):205, June 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869478.

Anonymous:1975:E

[302] Anonymous. Errata. Synthese, 31(1):207, June 1975. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869479;   http://link.
springer.com/content/pdf/10.1007/BF00869479.pdf.

Berlinski:1975:GEI

[303] David Berlinski. Guest editor’s introduction. Synthese, 31(2):209– 210, August 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485977.

Berlinski:1975:MMW

[304] David Berlinski. Mathematical models of the world. Synthese, 31(2): 211–227, August 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485978.

Sussmann:1975:CT

[305] Hector J. Sussmann. Catastrophe theory. Synthese, 31(2):229–270, August 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485979.

Intriligator:1975:AOC

[306] Michael D. Intriligator. Applications of optimal control theory in economics. Synthese, 31(2):271–288, August 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485980.

Gillespie:1975:PMM

[307] John V. Gillespie and Dina A. Zinnes. Progressions in mathematical models of international conflict. Synthese, 31(2):289–321, August 1975. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485981.

Samuelson:1975:MPA

[308] Paul A. Samuelson. Maximum principles in analytical economics. Synthese, 31(2):323–344, August 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485982.

Smale:1975:GAE

[309] Steve Smale. Global analysis and economics. Synthese, 31(2):345– 358, August 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485983.

Thom:1975:DMS

[310] R. Thom. D’un modèle de la science à une science des modèles. (French) [From a model of science to a science of models]. Synthese, 31(2):359– 374, August 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485984.

Anonymous:1975:BJR

[311] Anonymous. Books and journals received. Synthese, 31(2):375–376, August 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485985.

Anonymous:1975:Aa

[312] Anonymous. Announcement. Synthese, 31(2):377–379, August 1975. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485986.

Ware:1975:SBP

[313] Robert X. Ware. Some bits and pieces. Synthese, 31(3–4):379–393, September 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485210.

Cartwright:1975:SRA

[314] Helen M. Cartwright. Some remarks about mass nouns and plurality. Synthese, 31(3–4):395–410, September 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485211.

Laycock:1975:TM

[315] Henry Laycock. Theories of matter. Synthese, 31(3–4):411–442, September 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485212.

Cook:1975:UQ

[316] Kathleen C. Cook. On the usefulness of quantities. Synthese, 31(3–4): 443–457, September 1975. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485213.

Burge:1975:MTC

[317] Tyler Burge. Mass terms, count nouns, and change. Synthese, 31(3–4): 459–478, September 1975. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485214.

Grandy:1975:ST

[318] Richard E. Grandy. Stuff and things. Synthese, 31(3–4):479–485, September 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485215.

Chellas:1975:QQ

[319] Brian F. Chellas. Quantity and quantification. Synthese, 31(3–4):487– 491, September 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485216.

Bealer:1975:PM

[320] G. Bealer. Predication and matter. Synthese, 31(3–4):493–508, September 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485217.

Zemach:1975:ATO

[321] E. Zemach. On the adequacy of a type ontology. Synthese, 31(3–4): 509–515, September 1975. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485218.

Parsons:1975:AMT

[322] Terence Parsons. Afterthoughts on mass terms. Synthese, 31(3–4):517– 521, September 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485219.

Pelletier:1975:BRW

[323] Francis Jeffry Pelletier. A bibliography of recent work on mass terms. Synthese, 31(3–4):523–526, September 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485220.

Cohen:1975:A

[324] Prof. R. S. Cohen. Announcement. Synthese, 31(3–4):527–530, September 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485221.

Burian:1975:CCC

[325] Richard M. Burian. Conceptual change, cross-theoretical explanation, and the unity of science. Synthese, 32(1–2):1–28, March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485111.

Harris:1975:CT

[326] John H. Harris. On comparing theories. Synthese, 32(1–2):29–76, March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485112.

Redhead:1975:SIR

[327] M. L. G. Redhead. Symmetry in intertheory relations. Synthese, 32(1–2): 77–112, March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485113.

Watanabe:1975:NHD

[328] Satosi Watanabe. Needed: A historico-dynamical view of theory change. Synthese, 32(1–2):113–134, March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485114.

Spector:1975:RMR

[329] Marshall Spector. Russell’s maxim and reduction as replacement. Synthese, 32(1–2):135–176, March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485115.

Hooker:1975:PMP

[330] C. A. Hooker. Philosophy and meta-philosophy of science: Empiricism, Popperianism and realism. Synthese, 32(1–2):177–231, March 1975. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485116.

Cohen:1975:DHW

[331] Robert S. Cohen and Joseph Agassi. Dinosaurs and horses, or: Ways with nature. Synthese, 32(1–2):233–247, March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485117.

Anonymous:1975:CIU

[332] Anonymous. Communications of the international union of history and philosophy of science, division of logic, methodology, and philosophy of science. Synthese, 32(1–2):249–265, March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485118.

Anonymous:1975:IUH

[333] Anonymous. International Union of the History and Philosophy of Science, Division of Logic, Methodology and Philosophy of Science Bulletin No. 1. Synthese, 32(1–2):267–268, March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485119.

Anonymous:1975:Ab

[334] Anonymous. Announcement. Synthese, 32(1–2):269, March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485120.

Martin:1976:NLU

[335] Robert L. Martin. Are natural languages universal? Synthese, 32(3–4): 271–291, April 1976. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484779.

Jones:1976:GSS

[336] A. J. I. Jones. Generative semantics: Some test cases. Synthese, 32 (3–4):293–307, April 1976. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484780.

Needham:1976:SPV

[337] Paul Needham. The speaker’s point of view. Synthese, 32(3–4):309– 327, April 1976. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484781.

Richards:1976:ALP

[338] Barry Richards. Adverbs: From a logical point of view. Synthese, 32 (3–4):329–372, April 1976. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484782.

Steinberg:1976:CPP

[339] Danny D. Steinberg. Competence, performance and the psychological invalidity of Chomsky’s grammar. Synthese, 32(3–4):373–386, April 1976. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00484783.

Harman:1976:KC

[340] Gilbert Harman. Katz’ credo. Synthese, 32(3–4):387–394, April 1976. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00484784.

Grandy:1976:ALE

[341] Richard E. Grandy. Anadic logic and English. Synthese, 32(3–4):395– 402, April 1976. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484785.

Friedman:1976:UCS

[342] Joel Friedman. The universal class has a Spinozistic partitioning. Synthese, 32(3–4):403–418, April 1976. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484786.

Lake:1976:SPC

[343] John Lake. Spinozistic partitions of classes. Synthese, 32(3–4):419– 421, April 1976. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484787.

Anonymous:1976:BJR

[344] Anonymous. Books and journals received. Synthese, 32(3–4):423– 425, April 1976. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484788.

Anonymous:1976:A

[345] Anonymous. Announcement. Synthese, 32(3–4):427, April 1976. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00484789.

Anonymous:1976:E

[346] Anonymous. Errata. Synthese, 32(3–4):429, April 1976. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00484790;   http://link.
springer.com/content/pdf/10.1007/BF00484790.pdf.

Bellacicco:1976:FC

[347] Antonio Bellacicco. Fuzzy classifications. Synthese, 33(1):273–281, June 1976. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485447.

Albrecht-Buehler:1976:NEV

[348] G. Albrecht-Buehler. Numerical evaluation of the validity of experimental proofs in biology. Synthese, 33(1):283–312, June 1976. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485448.

Cherry:1976:EEH

[349] Christopher Cherry. Explanation and explanation by hypothesis. Synthese, 33(1):315–339, June 1976. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485449.

Freundlich:1976:RR

[350] Yehudah Freundlich. Resurrecting the ravens. Synthese, 33(1):341– 354, June 1976. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485450.

Sachs:1976:PR

[351] Mendel Sachs. Popper and reality. Synthese, 33(1):355–369, June 1976. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485451.

Elster:1976:NHS

[352] Jon Elster. A note on hysteresis in the social sciences. Synthese, 33(1): 371–391, June 1976. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485452.

Fishburn:1976:NP

[353] Peter C. Fishburn. Noncompensatory preferences. Synthese, 33(1):393– 403, June 1976. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485453.

Ballonoff:1976:ST

[354] Paul A. Ballonoff. Stray theory. Synthese, 33(1):405–418, June 1976. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485454.

Ramsay:1976:ARP

[355] J. O. Ramsay. Algebraic representation in the physical and behavioral sciences. Synthese, 33(1):419–453, June 1976. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485455.

Osborne:1976:UTD

[356] Dale K. Osborne. Unified theory of derived measurement. Synthese, 33 (1):455–481, June 1976. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485456.

Luce:1976:QER

[357] R. Duncan Luce and Louis Narens. A qualitative equivalent to the relativistic addition law for velocities. Synthese, 33(1):483–487, June 1976. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485457.

Demopoulos:1976:CN

[358] William Demopoulos. Critical notice. Synthese, 33(1):489–504, June 1976. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485458.

Craig:1976:SEF

[359] Edward Craig. Sensory experience and the foundations of knowledge. Synthese, 33(2–4):1–24, September 1976. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484707.

Pears:1976:CCP

[360] D. F. Pears. The causal conditions of perception. Synthese, 33(2–4):25– 40, September 1976. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484708.

Bradie:1976:CTP

[361] Michael P. Bradie. The causal theory of perception. Synthese, 33(2– 4):41–74, September 1976. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484709.

Clark:1976:OFL

[362] Romane Clark. Old foundations for a logic of perception. Synthese, 33(2– 4):75–99, September 1976. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00484710.

Sober:1976:MR

[363] Elliott Sober. Mental representations. Synthese, 33(2–4):101–148, September 1976. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484711.

Howell:1976:OPM

[364] Robert Howell. Ordinary pictures, mental representations, and logical forms. Synthese, 33(2–4):149–174, September 1976. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00484712.

Krausser:1976:KSC

[365] Peter Krausser. Kant’s schematism of the categories and the problem of pattern recognition. Synthese, 33(2–4):175–192, September 1976. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00484713.

Saford:1976:CSV

[366] David H. Saford. Competing semantics of vagueness: Many values versus super-truth. Synthese, 33(2–4):195–210, September 1976. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00484714.

Margalit:1976:VV

[367] Avishai Margalit. Vagueness in vogue. Synthese, 33(2–4):211–221, September 1976. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484715.

Mansfield:1976:BVS

[368] Richard Mansfield and John Dawson. Boolean-valued set theory and forcing. Synthese, 33(2–4):223–252, September 1976. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00484716.

Weiss:1976:SFW

[369] Stephen E. Weiss. The sorites fallacy: What difference does a peanut make? Synthese, 33(2–4):253–272, September 1976. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00484717.

Salmon:1977:P

[370] Wesley C. Salmon. Preface. Synthese, 34(1):ii–2, January 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485635.

Salmon:1977:PHR

[371] Wesley C. Salmon. The philosophy of Hans Reichenbach. Synthese, 34(1): 5–88, January 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485636.

Clendinnen:1977:IPT

[372] F. John Clendinnen. Inference, practice and theory. Synthese, 34(1): 89–132, January 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485637.

VanFraassen:1977:RF

[373] Bas C. Van Fraassen. Relative frequencies. Synthese, 34(2):133–166, February 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485800.

Rogers:1977:PTF

[374] Ben Rogers. The probabilities of theories as frequencies. Synthese, 34 (2):167–183, February 1977. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485801.

Fetzer:1977:RRC

[375] James H. Fetzer. Reichenbach, reference classes, and single case ‘probabilities’. Synthese, 34(2):185–217, February 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485802.

Glymour:1977:RE

[376] Clark Glymour. Reichenbach’s entanglements. Synthese, 34(2):219– 235, February 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485803.

Lehrer:1977:RC

[377] Keith Lehrer. Reichenbach on convention. Synthese, 34(2):237–248, February 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485804.

Kamlah:1977:HRR

[378] Andreas Kamlah. Hans Reichenbach’s relativity of geometry. Synthese, 34(3):249–263, March 1977. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485877.

Beauregard:1977:RC

[379] Laurent A. Beauregard. Reichenbach and conventionalism. Synthese, 34 (3):265–280, March 1977. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485878.

Grunbaum:1977:GRD

[380] Adolf Grünbaum and Allen I. Janis. The geometry of the rotating disk in the special theory of relativity. Synthese, 34(3):281–299, March 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485879.

Salmon:1977:CPA

[381] Merrilee H. Salmon. Consistency proofs for applied mathematics. Synthese, 34(3):301–312, March 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485880.

Nilson:1977:HRL

[382] Donald Richard Nilson. Hans Reichenbach on the logic of quantum mechanics. Synthese, 34(3):313–360, March 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485881.

Mackie:1977:DGC

[383] J. L. Mackie. Dispositions, grounds, and causes. Synthese, 34(4):361– 369, April 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485646.

Essler:1977:SWO

[384] Wilhelm K. Essler and Rainer Trapp. Some ways of operationally introducing dispositional predicates with regard to scientific and ordinary practice. Synthese, 34(4):371–396, April 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485647.

Fetzer:1977:WD

[385] James H. Fetzer. A world of dispositions. Synthese, 34(4):397–421, April 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485648.

Levi:1977:SDC

[386] Isaac Levi. Subjunctives, dispositions and chances. Synthese, 34(4): 423–455, April 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485649.

Tuomela:1977:DRE

[387] Raimo Tuomela. Dispositions, realism, and explanation. Synthese, 34 (4):457–478, April 1977. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485650.

Pastin:1977:CE

[388] Mark Pastin. Counterfactuals in epistemology. Synthese, 34(4):479– 495, April 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485651.

Feldman:1977:AW

[389] Fred Feldman. On the analysis of warranting. Synthese, 34(4):497– 512, April 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485652.

Anonymous:1977:Aa

[390] Anonymous. Announcement. Synthese, 34(4):513, April 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485653.

Hardegree:1977:RLQ

[391] Gary M. Hardegree. Reichenbach and the logic of quantum mechanics. Synthese, 35(1):3–40, March 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485434;   http://link.springer.com/content/
pdf/10.1007/BF00485434.pdf.

Jones:1977:CAC

[392] Roger Jones. Causal anomalies and the completeness of quantum theory. Synthese, 35(1):41–78, March 1977. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485435;   http://link.
springer.com/content/pdf/10.1007/BF00485435.pdf.

Grossman:1977:MIQ

[393] Neal Grossman. Metaphysical implications of the quantum theory. Synthese, 35(1):79–97, March 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485436;   http://link.springer.com/content/
pdf/10.1007/BF00485436.pdf.

McMahon:1977:GMT

[394] William E. McMahon. A generative model for translating from ordinary language into symbolic notation. Synthese, 35(1):99–116, March 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485437;   http:
//link.springer.com/content/pdf/10.1007/BF00485437.pdf.

Schuster:1977:ACR

[395] Cynthia Schuster. Appreciation and criticism of Reichenbach’s metaethics: Achilles’ heel of the system? Synthese, 35(1):117–126, March 1977.

CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485438;   http:
//link.springer.com/content/pdf/10.1007/BF00485438.pdf.

deBeauregard:1977:TLD

[396] O. Costa de Beauregard. Two lectures on the direction of time. Synthese, 35(2):129–154, June 1977. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485493.

Sklar:1977:WMR

[397] Lawrence Sklar. What might be right about the causal theory of time. Synthese, 35(2):155–171, June 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485494.

vonBretzel:1977:CPT

[398] Philip von Bretzel. Concerning a probabilistic theory of causation adequate for the causal theory of time. Synthese, 35(2):173–190, June 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485495.

Salmon:1977:LMC

[399] Wesley C. Salmon. Laws, modalities and counterfactuals. Synthese, 35 (2):191–229, June 1977. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485496.

Jobe:1977:RTN

[400] Evan K. Jobe. Reichenbach’s theory of nomological statements. Synthese, 35(2):231–254, June 1977. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485497.

Goldman:1977:PO

[401] Alvin I. Goldman. Perceptual objects. Synthese, 35(3):257–284, September 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485548.

Castaneda:1977:PBS

[402] Hector-Neri Castañeda. Perception, belief, and the structure of physical objects and consciousness. Synthese, 35(3):285–351, September 1977.

CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485549.

Farrell:1977:PEV

[403] B. A. Farrell. On the psychological explanation of visual perception. Synthese, 35(3):353–379, September 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485550.

Hall:1977:SN

[404] Richard J. Hall. Seeing and naming. Synthese, 35(3):381–393, September 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485551.

Suppes:1977:VSE

[405] Patrick Suppes. Is visual space Euclidean? Synthese, 35(4):397–421, December 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485624.

Battro:1977:VRS

[406] Antonio M. Battro. Visual Riemannian space versus cognitive Euclidean space. Synthese, 35(4):423–429, December 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485625.

Lofgren:1977:EEP

[407] Lars Löfgren. On existence and existential perception. Synthese, 35(4): 431–445, December 1977. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485626.

Peri:1977:PUP

[408] S. Peri. The problem of universals and its perceptual correlates. Synthese, 35(4):447–456, December 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485627.

Anonymous:1977:BJR

[409] Anonymous. Books and journals received. Synthese, 35(4):457–462, December 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485628.

Anonymous:1977:Ab

[410] Anonymous. Announcements. Synthese, 35(4):463–469, December 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485629.

Anonymous:1977:EI

[411] Anonymous. Editorial introduction. Synthese, 36(1):3, September 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485687.

Giere:1977:ABC

[412] Ronald N. Giere. Allan Birnbaum’s conception of statistical evidence. Synthese, 36(1):5–13, September 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485688.

Giere:1977:PAB

[413] Ronald N. Giere. Publications by Allan Birnbaum. Synthese, 36(1):15– 17, September 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485689.

Birnbaum:1977:NPT

[414] Allan Birnbaum. The Neyman–Pearson theory as decision theory, and as inference theory; with a criticism of the Lindley–Savage argument for Bayesian theory. Synthese, 36(1):19–49, September 1977. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485690.

Lindley:1977:DBI

[415] D. V. Lindley. The distinction between inference and decision. Synthese, 36(1):51–58, September 1977. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485691.

Pratt:1977:DSE

[416] John W. Pratt. Decisions as statistical evidence and Birnbaum’s ‘confidence concept’. Synthese, 36(1):59–69, September 1977. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485692.

Smith:1977:ABD

[417] Cedric A. B. Smith. The analogy between decision and inference. Synthese, 36(1):71–85, September 1977. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485693.

Kyburg:1977:DCU

[418] Henry E. Kyburg, Jr. Decisions, conclusions, and utilities. Synthese, 36 (1):87–96, September 1977. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485694.

Neyman:1977:FPF

[419] J. Neyman. Frequentist probability and frequentist statistics. Synthese, 36 (1):97–131, September 1977. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485695.

Lecam:1977:NME

[420] Lucien Lecam. A note on metastatistics or ‘an essay toward stating a problem in the doctrine of chances’. Synthese, 36(1):133–160, September 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485696.

Kiefer:1977:FST

[421] J. Kiefer. The foundations of statistics — are there any? Synthese, 36(1): 161–176, September 1977. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485697.

Anonymous:1977:E

[422] Anonymous. Errata. Synthese, 36(1):177, September 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485698;   http://link.
springer.com/content/pdf/10.1007/BF00485698.pdf.

Rosenkrantz:1977:S

[423] R. D. Rosenkrantz. Support. Synthese, 36(2):181–193, October 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00486112.

Martin-Lof:1977:ETC

[424] P. Martin-Löf. Exact tests, confidence regions and estimates. Synthese, 36 (2):195–206, October 1977. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00486113.

Benenson:1977:RFD

[425] F. C. Benenson. Randomness and the frequency definition of probability. Synthese, 36(2):207–233, October 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00486114.

Spielman:1977:PPB

[426] Stephen Spielman. Physical probability and Bayesian statistics. Synthese, 36(2):235–269, October 1977. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00486115.

Diaconis:1977:FFF

[427] Persi Diaconis. Finite forms of de Finetti’s theorem on exchangeability. Synthese, 36(2):271–281, October 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00486116.

Berlinski:1977:EI

[428] David Berlinski. Editor’s introduction. Synthese, 36(2):285–286, October 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00486117.

Fishburn:1977:MIP

[429] Peter C. Fishburn. Models of individual preference and choice. Synthese, 36(3):287–314, November 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00486023.

Thompson:1977:BTI

[430] John M. T. Thompson and Giles W. Hunt. A bifurcation theory for the instabilities of optimization and design. Synthese, 36(3):315–351, November 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00486024.

Nozick:1977:AM

[431] Robert Nozick. On Austrian methodology. Synthese, 36(3):353–392, November 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00486025.

Anonymous:1977:Ac

[432] Anonymous. Announcement. Synthese, 36(3):393–395, November 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00486026.

Salmon:1977:OHR

[433] Wesley C. Salmon. Objectively homogeneous reference classes. Synthese, 36(4):399–414, December 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00486104.

Humphreys:1977:RIH

[434] Paul W. Humphreys. Randomness, independence, and hypotheses. Synthese, 36(4):415–426, December 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00486105.

Suppes:1977:URR

[435] Patrick Suppes and Mario Zanotti. On using random relations to generate upper and lower probabilities. Synthese, 36(4):427–440, December 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00486106.

Havranek:1977:TMT

[436] Tomás Havránek. Towards a model theory of statistical theories. Synthese, 36(4):441–458, December 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00486107.

Bigelow:1977:SP

[437] John C. Bigelow. Semantics of probability. Synthese, 36(4):459–472, December 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00486108.

Laddaga:1977:LCP

[438] Robert Laddaga. Lehrer and the consensus proposal. Synthese, 36(4): 473–477, December 1977. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00486109. See reply [573].

Freudenthal:1977:R

[439] Hans Freudenthal. Review. Synthese, 36(4):479–492, December 1977. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00486110.

Kyburg:1977:RPF

[440] Henry E. Kyburg, Jr. Reply to Professor Freudenthal. Synthese, 36(4): 493–498, December 1977. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00486111.

Floistad:1978:MBS

[441] Guttorm Fløistad. Mind and body in Spinoza’s ethics. Synthese, 37(1): 1–13, January 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869439.

Jarrett:1978:LSS

[442] Charles Jarrett. The logical structure of Spinoza’s ethics, Part I. Synthese, 37(1):15–65, January 1978. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869440.

Friedman:1978:OSE

[443] Joel I. Friedman. An overview of Spinoza’s ehics. Synthese, 37(1):67– 106, January 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869441.

Martens:1978:SA

[444] Stanley C. Martens. Spinoza on attributes. Synthese, 37(1):107–111, January 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869442.

Anonymous:1978:Ea

[445] Anonymous. Errata. Synthese, 37(1):113–114, January 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869443;   http://link.
springer.com/content/pdf/10.1007/BF00869443.pdf.

Sussmann:1978:CTA

[446] Héctor J. Sussmann and Raphael S. Zahler. Catastrophe theory as applied to the social and biological sciences: A critique. Synthese, 37(2):117– 216, February 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869575.

Sankoff:1978:PLV

[447] David Sankoff. Probability and linguistic variation. Synthese, 37(2):217– 238, February 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869576.

Bauer:1978:RUG

[448] Christian Bauer. A reflection on universal grammars. Synthese, 37(2): 239–251, February 1978. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869577.

S:1978:E

[449] W. C. S. Erratum. Synthese, 37(2):253, February 1978. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869578;   http://link.
springer.com/content/pdf/10.1007/BF00869578.pdf.

Kuipers:1978:GCI

[450] Theo A. F. Kuipers. On the generalization of the continuum of inductive methods to universal hypotheses. Synthese, 37(3):255–284, March 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00873242.

Rossi:1978:AIL

[451] Robert Joseph Rossi. Application of inductive logic to the analysis of construct validity. Synthese, 37(3):285–319, March 1978. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00873243.

Nowak:1978:LEA

[452] Stefan Nowak. Logical and empirical assumptions of validity of inductions. Synthese, 37(3):321–349, March 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00873244.

Gardenfors:1978:LR

[453] Peter Gärdenfors. On the logic of relevance. Synthese, 37(3):351– 367, March 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00873245.

Runnalls:1978:DLS

[454] A. R. Runnalls. Difficulties of the Lindley–Savage argument. Synthese, 37(3):369–385, March 1978. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00873246.

Lindenberg:1978:BP

[455] Siegwart Lindenberg and Paul Oppenheim. The bargain principle. Synthese, 37(3):387–412, March 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00873247.

Currie:1978:PEE

[456] Gregory Currie. Popper’s evolutionary epistemology: A critique. Synthese, 37(3):413–431, March 1978. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00873248.

Tichy:1978:NTS

[457] Pavel Tichý. A new theory of subjunctive conditionals. Synthese, 37 (3):433–457, March 1978. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00873249.

Pastin:1978:WRR

[458] Mark Pastin. Warranting reconsidered: Response to Feldman. Synthese, 37(3):459–464, March 1978. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00873250.

Feldman:1978:FCA

[459] Fred Feldman. Final comments on the analysis of warranting. Synthese, 37(3):465–469, March 1978. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00873251.

Tranoy:1978:NFS

[460] Knut Erik Tranöy. Normative foundations of science. Synthese, 37(3): 471–477, March 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00873252.

Anonymous:1978:Aa

[461] Anonymous. Announcement. Synthese, 37(3):478, March 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00873253.

Levin:1978:IRS

[462] Michael E. Levin and Margarita R. Levin. The independence results of set theory: An informal exposition. Synthese, 38(1):1–34, May 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00484751.

Butrick:1978:ST

[463] Richard Butrick. Shoenfield on n-tuples. Synthese, 38(1):35–37, May 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00484752.

Cooper:1978:LLE

[464] William S. Cooper. The logico-linguistic evidence underlying Montague’s language descriptions. Synthese, 38(1):39–71, May 1978. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00484753.

Boer:1978:PNF

[465] Steven E. Boër. Proper names and formal semiotic. Synthese, 38(1): 73–112, May 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484754.

Stone:1978:TWT

[466] John David Stone. Treatments of the word ‘true’ in Montague grammar. Synthese, 38(1):113–125, May 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484755.

Rosemont:1978:GEL

[467] Henry Rosemont, Jr. Gathering evidence for linguistic innateness. Synthese, 38(1):127–148, May 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00484756.

Churchland:1978:FLL

[468] Patricia Smith Churchland. Fodor on language learning. Synthese, 38(1): 149–159, May 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00484757.

Wasow:1978:CN

[469] Thomas Wasow. Critical notices. Synthese, 38(1):161–167, May 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00484758.

Fenstad:1978:CN

[470] Jens Erik Fenstad. Critical notices. Synthese, 38(1):169–173, May 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00484759.

Tichy:1978:VR

[471] Pavel Tichý. Verisimilitude revisited. Synthese, 38(2):175–196, June 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00486149.

Miller:1978:DBC

[472] David Miller. The distance between constituents. Synthese, 38(2):197– 212, June 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00486150.

Tuomela:1978:TDV

[473] Raimo Tuomela. Theory-distance and verisimilitude. Synthese, 38(2): 213–246, June 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00486151.

Mott:1978:VMS

[474] Peter L. Mott. Verisimilitude by means of short theorems. Synthese, 38(2):247–273, June 1978. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00486152.

Krajewski:1978:ATF

[475] Władysław Krajewski. Approximative truth of fact-statements, laws, and theories. Synthese, 38(2):275–279, June 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00486153.

Niiniluoto:1978:TCR

[476] Ilkka Niiniluoto. Truthlikeness: Comments on recent discussion. Synthese, 38(2):281–329, June 1978. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00486154.

Anonymous:1978:Eb

[477] Anonymous. Erratum. Synthese, 38(2):331, June 1978. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00486155;   http://link.
springer.com/content/pdf/10.1007/BF00486155.pdf.

Bealer:1978:IF

[478] George Bealer. An inconsistency in functionalism. Synthese, 38(3): 333–372, July 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00486635.

Laing:1978:ASD

[479] Richard Laing. Anomalies of self-description. Synthese, 38(3):373– 387, July 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00486636.

Boden:1978:AIP

[480] Margaret A. Boden. Artificial intelligence and Piagetian theory. Synthese, 38(3):389–414, July 1978. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00486637.

Dennett:1978:WYC

[481] Daniel C. Dennett. Why you can’t make a computer that feels pain. Synthese, 38(3):415–456, July 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00486638.

Bennett:1978:DIM

[482] Michael Bennett. Demonstratives and indexicals in Montague grammar. Synthese, 39(1):1–80, September 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485407.

Wasow:1978:CCT

[483] Thomas Wasow. On constraining the class of transformational languages. Synthese, 39(1):81–104, September 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485408.

Nelson:1978:OOB

[484] R. J. Nelson. Objects of occasion beliefs. Synthese, 39(1):105–139, September 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485409.

Thomason:1978:CN

[485] Richmond H. Thomason. Critical notice. Synthese, 39(1):141–154, September 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485410.

Parsons:1978:CN

[486] Terence Parsons. Critical notice. Synthese, 39(1):155–164, September 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485411.

Domotor:1978:AJU

[487] Zoltan Domotor. Axiomatization of Jeffrey utilities. Synthese, 39(2):165– 210, October 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485073.

Tuomela:1978:SAD

[488] Raimo Tuomela. On the structuralist approach to the dynamics of theories. Synthese, 39(2):211–231, October 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485074.

Rantala:1978:ONL

[489] Veikko Rantala. The old and the new logic of metascience. Synthese, 39 (2):233–247, October 1978. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00485075.

Hunter:1978:CNP

[490] Daniel Hunter and Reed Richter. Counterfactuals and Newcomb’s paradox. Synthese, 39(2):249–261, October 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00485076.

Ullmann-Margalit:1978:IHE

[491] Edna Ullmann-Margalit. Invisible-hand explanations. Synthese, 39(2): 263–291, October 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485077.

Bruter:1978:TCS

[492] C. P. Bruter. The theory of catastrophes: Some epistemological aspects. Synthese, 39(2):293–315, October 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00485078.

Meyers:1978:IP

[493] Robert G. Meyers. Indeterminacy and positivism. Synthese, 39(2):317– 324, October 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485079.

Johnson:1978:CBC

[494] D. M. Johnson. Can belief be commanded? Synthese, 39(2):325– 334, October 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485080.

Anonymous:1978:Ec

[495] Anonymous. Errata. Synthese, 39(2):335, October 1978. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485081;   http://link.
springer.com/content/pdf/10.1007/BF00485081.pdf.

Nelson:1978:CPD

[496] R. J. Nelson. The competence-performance distinction in mental philosophy. Synthese, 39(3):337–381, November 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00869555.

Sober:1978:CC

[497] Elliott Sober. Computability and cognition. Synthese, 39(3):383–399, November 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869556.

Shanon:1978:GCH

[498] Benny Shanon. The genetic code and human language. Synthese, 39(3): 401–415, November 1978. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869557.

Maslov:1978:MDP

[499] S. Yu. Maslov. Macroevolution as deduction process. Synthese, 39(3): 417–434, November 1978. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869558.

Rozeboom:1978:LUR

[500] William W. Rozeboom. The logic of unboundedly reactive systems. Synthese, 39(3):435–530, November 1978. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00869559.

Anonymous:1978:Ab

[501] Anonymous. Announcement. Synthese, 39(3):531, November 1978. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869560.

Knuuttila:1979:P

[502] Simo Knuuttila. Preface. Synthese, 40(1):1–2, January 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413942.

Spade:1979:RRM

[503] Paul Vincent Spade. Recent research on medieval logic. Synthese, 40(1): 3–18, January 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413943.

Pinborg:1979:ECL

[504] Jan Pinborg. The English contribution to logic before Ockham. Synthese, 40(1):19–42, January 1979. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413944.

Ebbesen:1979:DMA

[505] Sten Ebbesen. The dead man is alive. Synthese, 40(1):43–70, January 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413945.

Dazeley:1979:IAL

[506] Howard L. Dazeley and Wolfgang L. Gombocz. Interpreting Anselm as logician. Synthese, 40(1):71–96, January 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413946.

Serene:1979:RGI

[507] Eileen F. Serene. Robert Grosseteste on induction and demonstrative science. Synthese, 40(1):97–115, January 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413947.

Murdoch:1979:PAF

[508] John E. Murdoch. Propositional analysis in fourteenth-century natural philosophy: A case study. Synthese, 40(1):117–146, January 1979. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413948.

Sylla:1979:PFN

[509] Edith Dudley Sylla. The a posteriori foundations of natural science. Synthese, 40(1):147–187, January 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413949.

Knuuttila:1979:CCF

[510] Simo Knuuttila and Anja Inkeri Lehtinen. Change and contradiction: A fourteenth-century controversy. Synthese, 40(1):189–207, January 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413950.

Anonymous:1979:Ea

[511] Anonymous. Errata. Synthese, 40(1):208, January 1979. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413951;   http://link.
springer.com/content/pdf/10.1007/BF00413951.pdf.

Matthews:1979:GSL

[512] Robert J. Matthews. Are the grammatical sentences of a language a recursive set? Synthese, 40(2):209–224, February 1979. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485677.

Bauerle:1979:TLN

[513] Rainer Bäuerle. Tense logics and natural language. Synthese, 40(2):225– 230, February 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485678.

Kuhn:1979:PT

[514] Steven Kuhn. The pragmatics of tense. Synthese, 40(2):231–263, February 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485679.

Burge:1979:FH

[515] Tyler Burge. Frege and the hierarchy. Synthese, 40(2):265–281, February 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00485680.

Wright:1979:SAR

[516] Crispin Wright. Strawson on anti-realism. Synthese, 40(2):283–299, February 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485681.

Peters:1979:TCF

[517] Stanley Peters. A truth-conditional formulation of Karttunen’s account of presupposition. Synthese, 40(2):301–316, February 1979. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00485682.

Heny:1979:RA

[518] Frank Heny. Review article. Synthese, 40(2):317–352, February 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485683.

vanBenthem:1979:CN

[519] J. F. A. K. van Benthem. Critical notice. Synthese, 40(2):353–373, February 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00485684.

Engdahl:1979:CN

[520] Elisabet Engdahl. Critical notice. Synthese, 40(2):375–387, February 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485685.

Needham:1979:CN

[521] Paul Needham. Critical notice. Synthese, 40(2):389–408, February 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485686.

Sklar:1979:PTC

[522] Lawrence Sklar. Probability as a theoretical concept. Synthese, 40(3): 409–414, March 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413412.

Kyburg:1979:TA

[523] Henry E. Kyburg. Tyche and Athena. Synthese, 40(3):415–438, March 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413413.

Giere:1979:PN

[524] Ronald N. Giere. Propensity and necessity. Synthese, 40(3):439– 451, March 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413414.

Fetzer:1979:SSO

[525] James H. Fetzer and Donald E. Nute. Syntax, semantics, and ontology: A probabilistic causal calculus. Synthese, 40(3):453–495, March 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413415.

Smokler:1979:SCP

[526] Howard Smokler. Single-case propensities, modality, and confirmation. Synthese, 40(3):497–506, March 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413416.

Anonymous:1979:BJR

[527] Anonymous. Books and journals received. Synthese, 40(3):507–511, March 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00413417.

Anonymous:1979:IUH

[528] Anonymous. International union of the history and philosophy of science. Synthese, 40(3):513–514, March 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413418.

Almeder:1979:PM

[529] Robert Almeder. Peirce on meaning. Synthese, 41(1):1–24, May 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869647.

Delaney:1979:PAM

[530] C. F. Delaney. Peirce’s account of mental activity. Synthese, 41(1): 25–36, May 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869648.

Haack:1979:FN

[531] Susan Haack. Fallibilism and necessity. Synthese, 41(1):37–63, May 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869649.

Morgan:1979:MAI

[532] Charles G. Morgan. Modality, analogy, and ideal experiments according to C. S. Peirce. Synthese, 41(1):65–83, May 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00869650.

Skagestad:1979:CPB

[533] Peter Skagestad. C. S. Peirce on biological evolution and scientific progress. Synthese, 41(1):85–114, May 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00869651.

Anonymous:1979:Aa

[534] Anonymous. Announcements. Synthese, 41(1):115–116, May 1979. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869652.

Unger:1979:TNO

[535] Peter Unger. There are no ordinary things. Synthese, 41(2):117–154, June 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00869568.

Wheeler:1979:W

[536] Samuel C. Wheeler. On that which is not. Synthese, 41(2):155–173, June 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00869569.

Sanford:1979:NOC

[537] David H. Sanford. Nostalgia for the ordinary: Comments on papers by Unger and Wheeler. Synthese, 41(2):175–184, June 1979. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00869570.

Hausman:1979:GPC

[538] Alan Hausman. Goodman’s perfect communities. Synthese, 41(2):185– 237, June 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869571.

Smith:1979:CEP

[539] David Woodruff Smith. The case of the exploding perception. Synthese, 41(2):239–269, June 1979. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869572.

Bacon:1979:LFP

[540] John Bacon. The logical form of perception sentences. Synthese, 41(2): 271–308, June 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869573.

Carleton:1979:PO

[541] Lawrence Richard Carleton. Perceptual objections. Synthese, 41(2): 309–320, June 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00869574.

Walley:1979:VMC

[542] Peter Walley and Terrence L. Fine. Varieties of modal (classificatory) and comparative probability. Synthese, 41(3):321–374, July 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869449.

Stemmer:1979:PP

[543] Nathan Stemmer. Projectible predicates. Synthese, 41(3):375–395, July 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic).

URL http://link.springer.com/article/10.1007/BF00869450.

Radner:1979:PT

[544] Michael Radner. Possible theories. Synthese, 41(3):397–415, July 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869451.

Moulines:1979:TNE

[545] C. Ulises Moulines. Theory-nets and the evolution of theories: The example of Newtonian mechanics. Synthese, 41(3):417–439, July 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869452.

Stromholm:1979:R

[546] Per Strømholm. Review. Synthese, 41(3):441–450, July 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869453.

Smart:1979:CN

[547] J. J. C. Smart. Critical notice. Synthese, 41(3):451–459, July 1979. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00869454.

Domotor:1979:R

[548] Zoltan Domotor. Review. Synthese, 41(3):461–463, July 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00869455.

Edwards:1979:MFQ

[549] David A. Edwards. The mathematical foundations of quantum mechanics. Synthese, 42(1):1–70, September 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413704.

Nordin:1979:DLQ

[550] Ingemar Nordin. Determinism and locality in quantum mechanics. Synthese, 42(1):71–90, September 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413705.

Stairs:1979:AFI

[551] Allen Stairs. On Arthur Fine’s interpretation of quantum mechanics. Synthese, 42(1):91–100, September 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413706.

Demopoulos:1979:BRP

[552] William Demopoulos. Boolean representations of physical magnitudes and locality. Synthese, 42(1):101–119, September 1979. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413707.

Healey:1979:QRN

[553] Richard Healey. Quantum realism: Naïveté is no excuse. Synthese, 42(1): 121–144, September 1979. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413708.

Fine:1979:HCF

[554] Arthur Fine. How to count frequencies: A primer for quantum realists. Synthese, 42(1):145–154, September 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.com/

article/10.1007/BF00413709.

vanFraassen:1979:HVM

[555] Bas C. van Fraassen. Hidden variables and the modal interpretation of quantum theory. Synthese, 42(1):155–165, September 1979. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413710.

Lenker:1979:CCT

[556] Terry D. Lenker. Carathéodory’s concept of temperature. Synthese, 42(1): 167–171, September 1979. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413711.

Shrader-Frechette:1979:HEM

[557] K. S. Shrader-Frechette. High-energy models and the ontological status of the quark. Synthese, 42(1):173–189, September 1979. CODEN SYN-TAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.

springer.com/article/10.1007/BF00413712.

Weingard:1979:SPA

[558] Robert Weingard. Some philosophical aspects of black holes. Synthese, 42 (1):191–219, September 1979. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413713.

Anonymous:1979:Eb

[559] Anonymous. Erratum. Synthese, 42(1):221, September 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00413714;   http://link.
springer.com/content/pdf/10.1007/BF00413714.pdf.

McKirahan:1979:DPM

[560] R. McKirahan. Diodorus and Prior and the master argument. Synthese, 42(2):223–253, October 1979. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00413805.

Schrader:1979:SSP

[561] David E. Schrader. A solution to the stone paradox. Synthese, 42(2): 255–264, October 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413806.

Read:1979:SRV

[562] Stephen Read. Self-reference and validity. Synthese, 42(2):265–274, October 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413807.

Hale:1979:SGD

[563] Bob Hale. Strawson, Geach and Dummett on singular terms and predicates. Synthese, 42(2):275–295, October 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00413808.

Tennant:1979:LGI

[564] Neil Tennant. Language games and intuitionism. Synthese, 42(2):297– 314, October 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413809.

Burge:1979:CN

[565] Tyler Burge. Critical notice. Synthese, 42(2):315–334, October 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00413810.

Currie:1979:LPM

[566] Gregory Currie. Lakatos’s philosophy of mathematics. Synthese, 42(2): 335–351, October 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00413811.

Elgin:1979:QDS

[567] Catherine Z. Elgin. Quine’s double standard: Indeterminacy and quantifying in. Synthese, 42(3):353–377, November 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://link.springer.

com/article/10.1007/BF00414123.

Glouberman:1979:SBT

[568] M. Glouberman. A stratified bundle theory. Synthese, 42(3):379–410, November 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00414124.

Finocchiaro:1979:IPH

[569] Maurice A. Finocchiaro. On the importance of philosophy for history of science: Studies in the logic of erudition. Synthese, 42(3):411–441, November 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00414125.

Musgrave:1979:PP

[570] Alan Musgrave. Problems with progress. Synthese, 42(3):443–464, November 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.1007/
BF00414126.

Furberg:1979:R

[571] Mats Furberg. Review. Synthese, 42(3):465–473, November 1979. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00414127.

Anonymous:1979:Ab

[572] Anonymous. Announcement. Synthese, 42(3):475, November 1979. CO-DEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00414128.

Lehrer:1983:PAI

[573] Keith Lehrer and Carl Wagner. Probability amalgamation and the independence issue: A reply to Laddaga. Synthese, 55(3):339–346, June 1983. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL

http://link.springer.com/article/10.1007/BF00485827. See [438].